A Complete Bibliography of Communication Design Quarterly Review

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

23 September 2025 Version 1.35

Title word cross-reference

#HPV [12].

//doi.org/10.26818/9780814214466 [245]. //doi.org/10.37514/TPC [306]. //doi.org/10.37514/TPC-B.2022.1367 [306]. //doi.org/10.4324/978 [364]. //doi.org/10.4324/978-1-003-27441-4 [364]. //intermezzo.enculturation.net/14 [307]. //intermezzo.enculturation.net/14stedman-et-al/index.html [307].

14-stedman-et-al [307]. 19 [346, 258]. 1st [364].

2011 [107]. 2012 [43]. 2013 [64]. 2014 [101]. 2015 [108]. 2017 [138].

3 [81].

4 [70, 364].

7 [65].

9 [83, 63, 82]. 978 [83, 70, 63, 81, 82, 65]. 978-0-262-01795-4 [70]. 978-0-321-82168-3 [81]. 978-1-003-27441-4 [364]. 978-1-4493-6761-9 [82]. 978-1-4819-6006-9 [63]. 978-1-933820-14-9 [83]. 978-1-933820-18-7 [65].

Aberrance [361]. ability [213]. absence [99]. Academia [345]. Academic [306, 358, 311, 275, 227]. academics [36]. Accented [332]. Acceptability [361]. access [84]. Accessability [42]. Accessibility [322, 340, 202, 200, 37, 80, 199]. Accessible [323, 322, 352, 179, 50]. accuracy [100]. achieve [107]. across [67]. Act [173]. action [40]. activist [206]. Activists [321, 320]. acts [255]. adaptive [28]. Address [294, 104]. addressing [155]. adopt [128]. adoption [250, 252]. adults [39]. advisory [193]. advocacy [62, 159]. advocating [185]. affinity [251]. Affordable [173]. African [332]. afterlife [318]. Agboka [256]. Age [327, 204, 257]. Agencies [172]. Agency [71]. Agnew [363]. AI [354, 327]. aids [168]. al/index.html [307]. Alabama [363, 172]. algorithm [248]. Algorithms [222]. Alliance [214]. allies [254]. Along [236, 250]. Alternate [270]. Alvarez [291]. Amanda [262]. Amazon [63]. Amber [350]. ambient [4]. American [105]. Amid [280]. Amityville [42]. Amplifying [303, 350]. Amy [349]. analog [10]. analog-like [10]. Analysis [345, 311, 353, 251, 186, 137, 253, 279]. analytics [56]. Analyzing [142, 99, 62]. Anatomy [101, 349]. Angela [234]. Angeli [247]. Ann [283, 115]. Anthropocene [240]. Anywhere [81]. API [216]. apparent [98]. appeal [95]. appearance [213]. application [90]. applications [158, 67]. applied [79, 54]. Applying [324]. Approach [287, 167, 92, 144, 145, 137, 128, 61, 56, 77, 163]. Approaches [281, 150, 6, 245]. Apps [362]. Arbor [115]. architects [49, 192]. architecture [16, 143, 51, 68, 218]. archives [255]. area [37]. art [231]. Articulating [356]. artifact [172]. Artificial [330]. Arts [164, 294]. Arts-based [294]. as/is [245]. Aspects [116, 84, 103]. assemblage [17]. assembling [54]. Assessing [100]. assessment [230, 154]. Asynchronous [312]. attention [95, 268]. attributes [10]. ATTW [13]. audience [111]. Audiences [285, 84, 77]. audit [259]. Augmented [281]. Augmenting [322]. Aural [323]. Authorship [343]. Automated [267]. Available [311]. awareness [116]. awesome [90]. Awful [255].

B [163, 257, 233]. B.2022.1367 [306]. badges [92]. Balancing [336, 350]. barriers [53]. Based [302, 313, 342, 214, 181, 16, 127, 124, 77, 294]. Bayesian [230]. Baywood [42]. be [276, 183]. Beck [316]. become [192]. behavior [124]. beliefs [161]. Benjamin [262]. Bennett [284]. Bernadette [233]. Better [288, 92, 47]. Between [287, 347, 119, 89, 213]. Beyond [300, 191, 316, 283]. Biased [332]. Big [132, 262, 6, 29, 8]. Binds [296]. Biodigital [326]. Black [293]. Black-Led [293]. Bloomsbury [227]. blowing [175]. Blueprints [120]. board [193]. Boca [163]. Bodies [282, 277, 228]. Body [349]. Boehnert [227]. Boiarsky [139]. Bonni [351]. Book [70, 71, 228, 72, 227, 43, 42, 49, 349, 197, 187, 50, 188, 198, 232, 148, 219, 149, 232, 349]. borders [208]. Boulder [139]. boundaries [62]. Brenda [257]. Bridgeford [232]. bridging [152]. Brief [328, 52]. Brooklyn [83, 50]. build [30, 49]. builder [87]. Building [341, 81, 268, 295, 340, 294, 47, 92, 30, 90]. Buley [65]. Burton [351]. Business [107, 310]. businesses [181].

C [246, 307, 315, 364, 114, 257, 272, 139]. CA [357, 120, 129]. Caldwell [357]. Call [302]. Callum [348]. Cambridge [64, 164]. campaign [239]. Campbell [365]. Campos [316]. Can [300, 174, 30, 49, 136, 106, 176]. Cancer [363]. cape [182]. captioned [148]. card [251]. Care [324, 173, 77, 155]. Cargile [306, 358]. Cargo [24]. Carie [350]. Case [139, 260, 202, 179, 125]. CDQ [224]. Cemetery [236]. Centered [163, 364, 57, 185, 21, 159]. Century [363, 234, 234]. CER [296, 299]. CEUX [338]. chair [14, 26]. challenge [21]. Challenges [338, 334]. Change [304, 205, 191, 22, 137, 208]. Changemaking [300]. changes [115]. Changing [73, 110, 4, 267, 76]. channel [100]. Chaos [121]. characteristic [47]. Characterizing [10]. Charting [93]. Chase [315]. Chicago [228, 114, 108, 214]. Chickens [34]. Children [305]. Chinese [210]. Choong [59]. Chris [257]. Christa [228]. Christopher [263]. cinematic [105]. Circulation [304]. Citation [345, 343]. Citational [274]. City [348]. claims [180]. classes [146]. Classroom [341, 329, 333, 335, 254, 316]. Clay [63]. Clearinghouse [306, 223]. Clearly [351]. Cleary [290]. climate [205, 22, 137]. closed [148]. closed-captioned [148]. Clouds [308]. CO [139, 57, 104]. co-creative [57]. co-designing [104]. Coalition [295]. Coalitional [293, 304]. Coalitions [294]. code [193]. Coder [354]. Codes [355]. Collaboration [280, 310, 168, 54, 272]. Collaborative [282, 310, 40, 88, 125, 142]. collateral [184]. collective [152]. collectivity [240]. College [315, 261]. Collins [223]. Colorado [306, 139, 223, 358, 256, 359]. Colton [241]. Come [352]. Commemoration [346]. comments [248]. commerce [173]. Communicate [344]. Communicating [206, 171, 54, 351, 247, 246, 137, 143, 19, 15, 138]. Communication [287, 280, 232, 323, 315, 42, 329, 352, 64, 360, 291, 234, 272, 89, 333, 361, 139, 319, 286, 3, 347, 356, 126, 25, 109, 45, 189, 152, 230, 195, 100, 260, 242, 239, 174, 175, 254, 39, 265, 69, 141, 12, 4, 215, 200, 165, 87, 145, 137, 99, 88, 46, 235, 108, 194, 28, 125, 105, 6, 61, 20, 93, 226, 37, 21, 238, 209, 7, 1, 9, 8, 236, 2, 102, 122, 155, 183, 85, 103, 199, 55, 91, 237, 290, 129, 238, 350, 317]. communications [57, 258]. communicative [40]. communicators [196, 190, 256]. Communities [300, 68, 271]. Community [292, 298, 341, 302, 342, 297, 322, 299, 301, 295, 338, 139, 335, 294, 214, 215, 117, 134]. Community-Based [302, 214]. Community-Driven [295]. Community-Engaged [298, 342, 299, 301, 338, 294]. Community/Classroom [335]. companies [96]. Company [324]. comparative [253, 56]. Comparing [143]. Competencies [352]. competition [105]. Complex [15, 171, 259, 177, 57, 206, 137, 88, 143, 142]. complexity [18, 22, 54, 19, 17]. Complicating [319]. Composing [308, 40]. Composition [315, 133, 262]. compromises [46, 20]. computing [4]. Concepts [295]. Conference [315]. Conflict [352]. conflicting [270]. congressional [239]. Connect [180]. connected [50]. Connecting [357]. connectivity [4]. consent [217]. considerations [116, 84, 102]. considering [243, 128, 122, 161]. Conspiracy [255]. constituent [133]. Constructing [289, 210]. construction [18]. consumerism [96]. Content [289, 238, 259, 60, 251, 144, 202, 16, 119, 250, 173, 84, 140, 364, 357, 238]. content-based [16]. Contested [169]. Context [346, 116, 177, 214, 57, 123, 102, 131, 140, 155]. contexts [214, 186, 174, 161, 33]. Contextual [184]. contextualization [77]. Continue [328]. Contract [342]. contracts [21]. contributed [154]. Control [336, 270]. Conversation [298, 271, 186, 80]. Conversations [316]. conveyance [140]. Cook [306, 358]. Coping [286]. Copyright [70]. core [32]. Cormick [246]. corporate [85]. Counter [346]. Course [342, 338, 260, 146, 193, 200, 55]. Courses [310, 135, 253]. coursework [194]. Courtney [307]. COVID [286, 346, 258]. COVID-19 [346, 258]. cow [236]. Craig [246]. Crane [163, 306, 358]. CRC [163, 59, 365]. create [152]. CreateSpace [63]. Creating [83, 120, 58, 84, 34]. creation [202]. creative [57]. crises [152]. crisis [100, 204, 258]. critical [62]. cropping [184]. Cross [175, 105, 260, 33, 58, 59]. Cross-cultural [175, 105, 33, 58, 59]. cross-disciplinary [33]. cross-organizational [33]. cross-program [260]. Crossing [208]. crossroads [46]. crowd [110]. crowd-sourced [110]. CSIRO [246]. Cultivating [193, 200, 334, 159]. cults [24]. Cultural [299, 102, 294, 175, 105, 155, 161, 33, 58, 59]. Culture [77, 148, 38, 58, 102, 318]. culture-sensitive [58]. curricula [195, 76]. curriculum [191]. Currie [348]. customization [213]. Cynthia [245].

D [307, 271, 233, 50, 223]. Danforth [307]. Dànielle [263]. Data [66, 309, 141, 326, 166, 4, 132, 136, 6, 106, 29, 184, 8, 48, 262, 348]. data-driven [136]. database [274]. David [271, 233]. Davidson [264]. DC [72]. deaf [179]. Death [346]. deception [261]. decolonial [243]. decoloniality [243]. Decolonizing [243, 299]. dedicated [328]. Deep [242, 322]. Default [314]. Defining [1]. deliberation [205, 206]. demands [191]. denialism [205]. Descriptions [353]. Design [25, 341, 83, 303, 163, 323, 281, 81, 339, 153, 360, 7, 334, 335, 319, 340, 350, 325, 353, 126, 24, 259, 109, 158, 100, 162, 146, 242, 57, 239, 220, 185, 174, 217, 39, 110, 265, 201, 135, 5, 215, 179, 30, 165, 58, 46, 235, 194, 95, 160, 105, 6, 61, 93, 151, 231, 53, 111, 1, 9, 56, 89, 50, 8, 124, 2, 59, 102, 122, 155, 156, 3, 103, 65, 159, 47, 91, 121, 227, 272]. design-led [57]. designed [181]. designer [182]. Designers [292, 4, 200, 79]. Designing [146, 288, 355, 69, 202, 293, 364, 299, 352, 168, 99, 344, 127, 106, 161, 214, 104, 50, 327]. designs [143]. desk [112]. Develop [320, 69]. Developers [280, 216]. Developing [312, 77]. Development [304, 312, 235, 11, 106, 183, 270]. Devices [326]. DeVoss [263]. diabetes [167]. diagnosing [63]. diagramming [251]. Diagrams [120]. Diasporic [332]. Diego [43]. difference [213]. different [143]. Digital [121, 158, 260, 70, 321, 320, 324, 319, 214, 239, 269, 104, 92, 40, 133, 38, 10, 318]. digital-like [10]. Dirty [291]. Disability [42, 62, 199]. disambiguation [172]. disaster [49]. disciplinary [33]. Discipline [330, 223]. Disciplines [332]. Disciplining [330]. discourse [214]. discourses [210]. discovering [205]. discussion [205]. disembodied [258]. disinformation [257]. Dislike [285]. Dislike-Minded [285]. distributed [268]. Diverse [303]. diversity [261]. divide [104]. DIY [167, 154]. DJs [215]. do [192]. documentation [220, 216]. Documents [311, 277]. doi.org [306, 364, 245]. Double [296]. Double-Binds [296]. doula [99]. doulas [99]. Dr. [328, 329, 331, 332]. Drawing [145]. Driven [295, 136]. Duke [284]. duplication [172]. during [100, 258]. Dynamic [18]. Dynamics [285].

e-commerce [173]. e-health [95]. Each [345]. earth [315, 210]. ease [11]. Eble [234, 359]. eBooks [85]. ecocene [227]. Ecology [284, 227]. Ed [130, 364]. edge [114]. Edinburgh [348]. Edited [257, 263, 42, 262, 264]. edition [82]. editor [219, 203, 165, 171, 211, 156, 199]. Editorial [212, 189, 128, 25, 177, 141, 150, 13, 224, 140]. Eds [307, 257, 348, 238, 316, 317, 256, 350, 359, 245]. Education [355, 200]. effect [47].

Effectively [344]. Effects [280]. efforts [254]. electronic [180]. Elephant [362]. Eli [271]. embed [277]. Embodied [282, 317]. Embracing [319]. emergence [60]. Emergency [172, 185, 247]. Emergent [301, 154]. Emerging [289, 113, 175, 93, 183]. Empathetic [111]. Empathic [334]. Empathy [341]. empirical [157]. Employer [353]. employing [62]. Empowerment [343]. enabled [118]. enculturation [307]. enculturation/Intermezzo [307]. end [27]. energy [206]. Enfranchisement [360]. Engage [302]. Engaged [292, 298, 342, 299, 301, 338, 294]. Engagement [355, 193]. Engaging [162, 17]. Engineering [341, 233, 334, 139, 351]. engines [222]. English [332]. enhance [135, 56, 176]. enough [181]. enrich [174]. ensure [235]. ensuring [43]. Environmental [355, 317, 242, 204, 20, 240, 209, 208]. Environments [322, 92, 48]. equipment [166]. Equipping [256, 256]. Equitable [343]. equity [294]. Erin [359]. Establishing [354]. Estee [316]. ethic [154]. Ethical [261, 281, 268]. Ethics [324, 361]. Eugenic [363]. Euphemism [347]. Evaluating [355, 85]. evaluations [95]. Everyday [83, 178, 291]. everyone [50]. everything [115]. Everywhere [81]. evidence [127, 255]. evidence-based [127]. Evolving [190]. examine [74]. Examining [123, 332, 112, 176]. examples [48]. Exceptions [305]. exchange [89]. exchanges [100]. execution [91]. exemplar [145]. Exhibit [313]. Exhibit-Based [313]. expanding [151]. Expansion [361]. Expectations [353, 32]. Experience [259, 306, 337, 301, 358, 338, 362, 365, 152, 158, 251, 254, 150, 49, 151, 250, 218, 51, 102, 156, 183, 161, 65, 43]. Experiences [83, 120, 30, 50]. Expertise [101, 196]. Explicit [311]. Exploratory [164]. Exploring [287, 185, 37, 186]. extent [74].

F [234, 359]. fabric [269]. Facebook [118, 119]. facilitators [194]. facing [235]. Faris [307]. Farman [130]. feedback [92]. feminism [98]. feminized [134]. FemTech [326]. Fernweh [309]. Ferrara [83]. field [93, 231, 245]. fields [233, 34, 233]. film [105]. filtering [88, 68]. Finding [293]. firefighter [166]. First [234, 95, 91]. fitness [170]. Fitter [363]. five [146]. fix [55]. fixing [63]. FL [163]. Flesh [101]. Floor [326]. flow [63]. flu [154]. flux [228]. form [145, 131]. forms [67]. Fort [223]. Forum [110]. forums [205]. Foundational [299]. Fountain [101]. four [220]. fracturing [207]. Framework [160]. Frameworks [234, 243]. Framing [118]. Friend [328]. Frost [359]. functions [67]. Fundamentals [163]. future [189, 27, 5, 226, 171, 157, 8]. futures [270].

G [256]. gadfly [201]. gain [178].

Gallagher [318]. Gallaudet [72]. Gallon [129]. Game [83, 280, 220, 183]. Gaps [287, 201]. GenAI [314]. Gendered [359]. generation [9]. Generative [354, 327]. genres [249, 276]. geometric [145]. geovisualizations [86]. Gertz [244]. gesture [119]. get [55]. Getto [238]. Ghemawat [107]. Ginny [80]. git [270]. gizmos [166]. Glascott [257]. glass [94]. Glimpse [330]. Global [107, 225, 161, 50]. globally [69]. Godwin [256]. going [191]. Goldblatt [271]. Gonzalez [351]. Good [328, 181, 324]. Gore [349]. Governance [121]. Government [139]. GPF [145]. Grading [342]. Graduate [296, 288]. Graham [351, 245]. Graphics [346, 34]. Grassroots [321, 208]. Gray [285, 249]. Gross [101]. group [208]. groups [118]. Guest [203, 165, 141, 150, 171, 199]. Guide [120, 233, 63, 226, 65]. guidelines [43].

Guiseppe [238]. Gustav [364]. gynecology [98].

H [138, 58]. Haas [234]. hacking [240]. Halcyon [328, 329, 331, 332]. Happier [363]. harm [214]. Hart [264]. Hart-Davidson [264]. Hartson [43]. Harvard [107]. hashtag [67]. HDR [106]. Health [303, 286, 365, 152, 158, 175, 39, 153, 95, 20, 93, 96, 169, 98, 173, 84, 77, 155, 180, 170, 245]. Healthcare [287, 151, 127]. hearing [168]. Heat [344]. Hector [70]. hedonism [249]. help [110, 112, 47]. Helping [144]. Hesford [305]. heuristic [259, 61, 226, 155]. heuristics [150]. Hidden [297]. higher [200]. hijacking [175]. Hills [357, 129]. hip [215]. histories [270]. History [349]. holisticism [170]. Holmes [241]. home [252]. Honoring [329]. hop [215]. Horton [50]. hospitalist [97]. hospitals [97]. Howard [364]. http [307]. https [306, 364, 245]. Hubert [64]. Human [305, 186, 168, 21, 106]. human-information [21]. human-machine [168]. human-technology [186]. humane [6]. Humanitarian [305]. humanities [158, 260, 51, 164]. Humility [299]. Hutchinson [316]. Hybrid [240, 251, 61]. hydraulic [207].

Icon [39]. ideas [102]. Identifying [32, 76, 149]. Identities [296]. Identity [293, 172]. ideographic [239]. ideological [103]. IEEE [233]. illness [118]. image [19]. imagining [215]. Immigrant [291]. Impact [314]. impacts [27]. Impatient [167]. Implementing [235, 81]. Implications [332]. importance [93, 182]. impressions [95]. improve [39]. Inc. [58, 233]. Inclusive [325, 179]. Inclusivity [343]. Incomplete [237]. increasingly [33]. independent [220]. index.html [307].

India [104]. Indigenous [349, 104]. induction [230]. industrial [89]. Industry [345, 36]. inequity [276]. influenced [208]. inform [147]. Informal [355]. informating [166]. Information [323, 115, 171, 192, 365, 264, 24, 100, 63, 276, 104, 39, 110, 16, 5, 108, 143, 105, 79, 123, 21, 56, 89, 124, 34, 142, 15]. informing [97]. InfoVis [143]. Infrastructural [360, 266, 278, 274]. infrastructure [276, 269, 268, 273, 279, 161, 277, 270]. infrastructures [265, 275, 267]. injustice [229]. Innovative [306, 358, 149]. Insights [334, 178]. Instructor [288, 310]. integrated [45]. integrating [195, 102]. Integration [341]. Intelligence [330]. interaction [135, 21, 56]. interactions [186]. Interactive [340, 10, 86, 182]. Intercoder [354]. interdependent [202]. Interdisciplinary [309]. interface [201, 11]. Interfaces [83]. Intermezzo [307]. intermezzo.enculturation.net [307]. International [287, 69, 106, 84, 77, 155]. Internetworked [308]. Interrogating [359]. Intersection [42]. intersectional [202]. Intertwingled [115]. intervene [169]. intervention [274, 98]. Interventions [288, 152]. Intimate [326]. Introducing [313]. Introduction [15, 298, 265, 273, 102, 203, 165, 93, 123, 171, 211, 156, 199]. Inventing [308]. Investigating [258]. Invisibility [291]. invisible [31]. ISBN [83, 70, 81, 82]. Issue [298, 93, 102]. issues [88]. item [88]. item-to-item [88].

J [307, 315, 257, 120, 272, 255, 130, 241, 318, 245]. Jack [238]. James [315]. Jane [284]. Janice [80]. Jared [241]. Jason [237, 364, 272]. Jenell [245]. Jenny [255]. Jeremy [348]. Jessica [365]. Jewish [236]. Jim [264]. Joanna [227]. Job [353, 190]. John [83, 357, 308, 233, 318, 245]. Johnson [245]. Jolliffe [271]. Jonathan [285]. journals [275]. Journey [52]. Journeys [120]. judgments [181]. Juli [257]. Just [343]. Justice [348, 256, 33]. Justin [257].

Kadlec [81]. Kalbach [120]. Kate [306, 358]. Kaufmann [43, 50]. Kelli [306, 358]. Kenny [101]. Key [234, 95, 234]. King [350]. Kmiec [233]. Knoblauch [64]. Knowledge [64, 61, 356, 30, 117]. knowledges [40]. Knox [348]. Kyle [307].

L [121, 363, 49, 316, 350, 247, 245, 65]. Lab [101]. Labor [342, 297, 134, 291]. Labor-Based [342]. laboratory [252]. Labriola [238]. Laguna [357, 129]. Lancaster [350]. landowner [207]. landscape [110]. Language [129, 72]. Large [335]. Large-Scale [335]. Lawrence [328, 329, 330, 331, 332]. Layering [322]. Learning [271, 310, 162, 28, 105, 48, 248]. Led [293, 57, 254]. Legacy [329, 330, 331]. legal [21]. LEMAs [172]. Lens [332]. Lessons [325]. Letter [328]. Letterforms [263]. Letting [57]. Level [338, 162, 22, 86, 182, 147, 225]. Lewis [257]. Liberatory [339]. library [47]. Licastro [262]. lies [94]. Life [320, 324, 319, 269, 182, 178]. lifecycle [250]. Light [94]. like [10]. Limits [324]. Lines [347]. lip [201]. Lisa [42, 245]. Lisbon [41]. Listening [345, 225]. literacies [146, 133]. Literacy [321, 320, 323, 322, 319, 326, 327, 193, 173, 257, 271]. Littlefield [244]. living [111, 252]. Liza [218]. LMS [135]. Local [172, 58, 86, 161]. localization [182]. Localizing [342, 137]. locally [69]. Locative [130, 231]. Lockhart [257]. Logie [308]. Lois [363]. Longo [233]. lore [141]. Luen [59]. Lynch [245]. Lynn [365].

M [307, 234, 317]. M.I.S.E.R [100]. MA [164, 50]. machine [35, 168, 248]. Makerspace [283]. Making [296, 283, 272, 147, 277]. Management [289, 172, 185, 149, 194, 19, 108]. managers [134]. Managing [134, 121]. Manual [360]. Mapping [177, 322, 344, 242, 52, 225, 161, 120]. Maps [321, 204, 154]. marginalized [274]. margins [104]. marketing [250]. marketplace [16]. Mary [317]. Massachusetts [349]. Masters [338]. Masters-Level [338]. Material [138]. Materiality [349]. Materials [355, 261]. Matter [284]. Matters [263, 316]. Matthew [82]. McGreggor [348]. McNair [283]. me [170]. measuring [119]. Media [121, 83, 120, 285, 65, 116, 109, 27, 67, 148, 210, 113, 49, 125, 226, 231, 103, 50]. mediascape [125]. mediated [266]. medical [228, 62, 84, 77, 155, 247]. medicine [93, 127, 245]. Meloncon [42, 245]. memory [108]. messaging [217]. Method [313, 362, 251]. Methodological [336, 20]. methodologies [31, 151, 256]. Methodology [356, 98, 183]. Methods [345, 230, 228, 57, 128, 184, 178, 223]. MI [115]. Michael [218, 307]. Michelle [234, 359]. Michigan [283]. microinteractions [79]. middleware [158]. Miller [262]. Minded [285]. mindful [235]. mindfulness [152]. Mining [82]. mis [243]. Miscommunication [139]. misconduct [254]. misinformation [257]. mixed [128]. mixed-methods [128]. MMU [274]. Mobile [130, 44, 161]. Mobility [138]. model [126, 97, 16, 235]. modeling [248]. Models [312, 18]. modern [116]. modest [37]. Montfort [164]. Morgan [43, 348, 50]. Morse [72]. Morville [115]. Movement [70]. MRIs [34]. Mueller [223]. multi [137]. multi-sectoral [137]. Multidisciplinary [341, 92]. multilingual [202]. Multimodal [186, 146]. multimodality [136]. multiparty [186]. multiply [274]. must [276]. my [166].

N [244, 164, 223]. name [51]. Narrative [130, 92]. Narratives [303]. nature [122, 277]. Navigating [296, 33]. NCTE [315]. Need [292, 31, 128]. Needs [350, 104, 127]. negotiating [228, 181, 196]. negotiation [160]. Neoliberal [288]. net [127]. Network [223]. Networked [40]. Networks [271, 109, 88, 114]. next [9]. Nicole [263]. Nihilism [244]. Noble [222]. Nolen [244]. nonprofits [144]. nonrational [209]. normalcy [62]. Notes [14, 26]. nuanced [205]. NY [83, 49, 101, 58, 50, 130]. NYU [222].

Obamacare [173]. observation [216]. occupation [191]. Official [229]. oh [166]. Ohio [305, 291, 255, 245]. older [39]. one [207, 192, 65]. Online [109, 303, 310, 162, 146, 205, 69, 135, 133, 118, 136, 253, 127, 173, 68, 248]. Open [36, 25, 183]. Openness [336, 161]. opioid [214]. Opportunities [338, 260, 79]. oppression [221, 169, 222]. Ops [264]. O’Reilly [120, 82]. organizational [33]. Organizations [293, 63, 139]. organizing [54]. Other [345]. our [27, 74, 128]. outbreak [175]. Outcomes [252]. Outreach [355]. overview [52]. own [154]. Oxford [58]. oysters [277].

P [363, 107, 115]. page [145]. Pandemic [280]. paradigm [127]. Paradigms [289]. Pardha [43]. Parlor [263, 257, 308, 218]. Parrish [257]. part [144]. participants [162]. Participation [282, 267, 49]. Participatory [53, 178, 286, 98, 154]. Partnership [335]. Partnerships [297]. past [157]. path [236]. Pathologies [359]. pathways [190]. Patient [151, 100, 150]. patients [167, 97, 180]. Patrick [59]. patterns [12]. Pedagogical [195, 332]. pedagogies [256, 256]. Pedagogy [336, 301, 338, 334, 333, 335, 260, 257, 274, 147]. Pei [59]. Pei-Luen [59]. Pelvic [326]. People [357, 6]. Perception [355]. Perceptions [310, 326, 261]. personal [166].

Personalized [87]. personas [69, 75]. perspective [207]. perspectives [189, 220]. Persuasive [138, 87]. Peters [363]. Pflugfelder [138]. Pharmaceutical [96]. PhD [365]. Photography [362]. pigs [203]. pilot [113]. pitfalls [132]. Pittsburg [262]. Pittsburgh [271, 264]. place [176]. planning [145]. plans [171]. Plants [300]. platform [269]. Plato [203]. play [61]. Playful [83]. playlists [215]. Please [328]. Plocher [59]. Policy [300, 311, 54, 208]. Political [239, 314, 284, 208]. Politics [349, 227]. poorly [181]. popular [148]. populations [104]. Porn [68]. pornography [68]. positioning [274]. Possibilities [301, 205, 53]. Post [320, 295]. Post-Digital [320]. Post-Roe [295]. poster [38]. Postigo [70]. postindustrial [267]. potential [37, 266]. Potts [49]. power [178]. PowerPoint [64]. pp. [42]. Practical [163]. Practice [290, 162, 174, 196, 147, 306, 358]. practice-level [162, 147]. Practices [296, 321, 280, 343, 333, 130, 146, 274, 136, 125, 176]. Practicum [232]. pre [239]. pre-digital [239]. Prefab [11]. Preparing [194, 189, 76]. presence [99]. present [157]. presentation [87]. preservation [72, 315]. preserver [182]. Press [70, 306, 107, 114, 64, 58, 108, 358, 218, 256, 359, 241, 245, 363, 237, 228, 72, 232, 163, 263, 283, 315, 357, 257, 308, 262, 348, 305, 271, 291, 349, 284, 129, 316, 164, 255, 139, 59, 350, 318, 351, 365, 264, 222]. preventing [254]. prevention [254, 253]. primer [194]. Principles [174, 32, 153, 95]. prioritization [100]. Privacy [327, 316]. privileging [275]. Problem [88, 136, 11]. Problems [272]. Procedure [311]. Proceedings [171, 15]. Process [43, 105]. processing [48]. productive [86]. products [69, 59]. Produsers [27]. Profession [290]. Professional [232, 333, 361, 312, 347, 226, 199]. professionalization [134]. professionals [36, 189]. Program [341, 260, 99, 28]. Programmatic [288]. Programming [164]. programs [226, 55]. Project [335, 194, 286]. projects [235, 98]. promises [132]. promote [146]. Promoting [179, 99]. proposal [37]. proposing [16]. proposition [126]. prosperity [107]. prosumer [16]. protective [166]. protocol [100]. prototype [77]. prototype-based [77]. prototypes [176]. Prototyping [231]. providers [127]. Public [293, 286, 152, 158, 57, 99, 235, 231, 208]. public-facing [235]. Publicly [311]. Publishing [311, 246]. Pushing [62]. PwrdBy [214]. Pyla [43].

QR [355]. quality [43]. Quantifiable [170]. Queering [217]. Quesenbery [50]. Questioning [288]. questions [25, 36, 27, 20, 29]. Quick [355]. Quick-Response [355].

R [129, 256, 318]. Racial [337, 169]. racism [222]. Radical [361, 274]. rationale [145]. Raton [163]. Rau [59]. Re [122, 327, 282, 74, 128, 156]. Re-considering [122, 128]. re-examine [74]. re-thinking [156]. reactions [176]. ReaderCentric [16]. Reading [148, 133, 201]. real [48]. real-time [48]. Reality [281, 111]. really [75]. rearticulates [51]. Rebecca [256]. recommended [60]. Reconceptualizing [319]. records [180]. Recovery [214]. recruitment [261]. redesign [193]. Redish [80]. Redrawing [321]. Reducing [214]. Reflecting [156]. Reflexes [176]. regulated [174]. reimagine [200, 279]. reimagining [98, 199]. reinforce [222]. reintegrating [133]. Relational [283]. relationship [119]. relevance [85]. Reliability [354]. Religious [72]. Remediation [48]. Remote [280]. Report [338, 259, 251, 254, 250, 106]. Research [298, 345, 309, 282, 313, 299, 362, 365, 356, 294, 25, 36, 27, 141, 30, 31, 128, 136, 28, 79, 78, 37, 184, 157, 50, 65]. Researchers [292]. Researching [19, 147]. Resistance [71, 274]. Resounding [331]. Resource [302]. Resources [345, 60, 127]. Responding [152, 71]. Response [355, 49, 229]. responsibilities [210]. responsible [149]. Responsive [191, 81, 185]. results [36]. rethinking [194]. revealing [175]. revenge [203]. Review [121, 363, 83, 237, 63, 232, 246, 163, 307, 306, 290, 263, 43, 283, 107, 315, 364, 357, 114, 42, 257, 308, 81, 138, 262, 244, 348, 120, 49, 64, 101, 58, 305, 108, 271, 291, 349, 284, 233, 234, 129, 285, 358, 272, 238, 115, 218, 316, 317, 164, 50, 255, 139, 59, 130, 256, 350, 359, 241, 318, 351, 365, 82, 247, 245, 65, 264, 222, 223, 70, 71, 228, 72, 227, 148, 219, 149, 197, 187, 188, 198]. reVITALize [98]. reward [92]. Rewriting [253]. Rex [43]. Rhet [264]. Rhetoric [207, 315, 138, 264, 278, 72, 8, 255, 159, 363, 101, 218, 241, 245]. Rhetorical [249, 67, 175, 253, 108, 247, 42]. Rhetoricity [263]. Rhetorics [288, 93, 283, 305]. RHM [150]. Rice [255]. Riders [81]. Ridolfo [264]. Right [348]. Rights [70, 305]. rise [22, 86, 182, 225]. Risk [344, 209, 317, 230, 207, 113, 240, 154, 139]. risks [175, 210, 169, 225]. Robot [354]. Rock [204]. Roe [295]. Rogue [360]. Role [237, 70, 360, 267]. roles [206]. Room [362]. Rosenfeld [121, 83, 50, 65]. Rouge [317]. Routledge [364, 138, 101, 247, 290, 49, 272, 238, 317, 130]. Rowman [244]. rules [21]. Running [324]. Rural [271, 104]. Russell [82]. Ruszkiewicz [238]. Ryan [245].

S [307, 305, 108, 317, 50, 350, 241, 245, 222, 177]. S.S [245]. safety [127]. Salt [315]. Salvo [218]. Samuel [317]. San [43]. Sanchez [315]. Saving [119]. Scale [335]. SCCI [171]. scholar [274]. scholars [113]. scholarship [243]. Science [139, 356, 246, 235, 54, 246]. Scientific [356, 228, 137, 226, 34]. Scissors [324]. Scoping [309]. Scott [263, 245]. screen [180]. screenshot [184]. script [96]. sea [119, 22, 86, 182, 225]. search [222]. Sebastopol [120]. Second [298, 354, 82, 91]. Section [328, 123]. sectoral [137]. see [12]. Seeking [294, 204]. selection [259, 124]. selfies [209]. Semantic [115, 353]. sense [223]. sensitive [58]. Sensors [166]. SEO [195]. seriously [78]. services [247, 59]. sets [190]. settings [133]. sexual [217, 254, 253]. Shades [205]. Shared [310]. sharing [178]. Sharon [351]. Sheryl [238]. Shivers [283]. Shivers-McNair [283]. should [74]. SIGDOC [13, 229, 224]. sign [72]. Signs [72]. silence [204]. Sino [105]. Siri [330, 332]. Site [172, 274]. Sites [81, 276, 169]. situated [6]. Situating [282]. situation [44]. situational [279]. skill [190]. skills [73, 134]. small [181]. smart [168, 252]. Social [303, 226, 82, 116, 109, 27, 67, 210, 92, 113, 112, 256, 33, 103, 134, 49]. Socially [343, 119]. Society [64]. software [270]. solutions [104, 149]. Solving [272, 88]. Sons [233]. sorting [251, 248]. sorting-affinity [251]. Sound [323]. sounds [148]. Soundwriting [307]. source [124]. sourced [110]. Space [296]. spaces [61, 308]. speak [57]. speaks [94]. Special [328, 93, 123, 102]. Spinuzzi [63, 114]. Stakeholders [294]. Standing [204]. stasis [54]. State [237, 232, 305, 291, 234, 316, 255, 241, 318, 245, 210]. statement [229]. States [253]. Stedman [307]. Steve [241]. Still [163]. Stinson [317]. Stories [301, 333]. Story [302, 225, 130]. Storytelling [337]. Strain [363]. strategies [195, 135, 145]. Strategy [357, 301, 238, 91, 60, 144]. stream [12]. streamlining [259]. Street [360, 225]. structure [125]. Structured [289]. struggles [162, 147]. Student [296, 254, 310, 27, 261, 40]. Students [287, 335, 194, 76]. Studies [139, 186, 62, 231, 159, 42]. Studios [115, 220]. study [113, 202, 108, 125, 216, 252]. studying [63]. Subverting [70]. success [76]. Sun [58]. SUNY [350]. supplement [135]. Support [237, 303, 288, 295, 118, 112, 266, 33]. supremacy [315]. surveillance [316, 154]. survey [36]. Surveying [280]. survival [65]. Sustainable [300]. sustained [268]. Sustaining [297]. Swarts [237]. SWOT [137]. Symposium [15, 171]. system [100, 18, 276]. Systems [340, 10, 56]. Szuc [50].

T [72, 232, 257, 101, 238, 350]. Tables [352]. tactic [175]. Tactical [320, 360]. tagging [68]. Taking [78]. talking [12]. Tara [257]. tasks [142]. Taste [285]. TC [76]. Teach [300, 251]. Teaching [232, 337, 339, 234, 74, 179, 91, 232]. team [65]. teams [268, 54]. Technical [41, 287, 237, 232, 290, 323, 293, 42, 329, 352, 101, 360, 125, 233, 234, 250, 129, 272, 238, 317, 333, 361, 350, 347, 356, 126, 45, 189, 195, 260, 239, 174, 254, 202, 141, 200, 112, 108, 28, 196, 226, 123, 37, 190, 236, 256, 183, 103, 199, 55, 91]. Technological [320]. Technologies [332, 167, 150, 165, 168, 169, 130]. Technology [70, 71, 138, 46, 139, 214, 186, 58, 250, 241, 161, 252, 365, 244]. technology-based [214]. tell [136, 106]. Telling [5]. Testing [86, 213, 340, 221, 179]. Teston [228]. text [275]. text-privileging [275]. Tham [364, 272]. Tharon [364]. Their [287]. themselves [277]. Theoretical [234, 150]. Theories [256, 245]. Theorizing [201]. Theory [196, 278, 255, 54]. thing [276]. things [78, 284]. Thinking [272, 335, 156]. three [144]. three-part [144]. thyroid [38]. Tim [81]. Time [296, 48]. times [73]. today [189]. Tom [59]. Tone [357]. Tool [309, 304, 112]. Tools [354, 344, 214, 206, 21]. topic [248]. topics [76]. Topsight [63]. TPC [337, 336, 243, 295]. TPC-B.2022.1367 [306]. TPC-UX [336]. Tracing [38, 304]. trackers [170]. Trained [101]. Training [326, 252]. transactional [235]. transdisciplinary [54]. transformation [125]. Transforming [21]. Translation [356, 35]. Transnational [320]. Transparent [311]. Transportation [138]. trauma [100]. trope [170]. Tuning [307]. Twentieth [363]. Twentieth-Century [363]. Twenty [234]. Twenty-First [234]. Twilight [325]. Twitter [12, 112]. two [68, 48]. Type [263].

U [222]. U.S. [353]. ubiquitous [4]. UI [353]. Uncertain [237]. uncertainty [22, 228]. undergraduate [162]. underrepresented [274]. Understanding [92, 79, 353, 111]. United [253]. Universities [287]. University [363, 237, 228, 72, 232, 306, 283, 114, 262, 348, 64, 58, 305, 108, 271, 291, 349, 284, 358, 316, 255, 139, 256, 359, 241, 318, 245, 264, 258, 234]. Unlikely [254]. unpredictable [247]. Update [318]. Updates [224]. Urban [271]. Usability [221, 365, 126, 167, 162, 186, 185, 179, 151, 80, 86, 23, 155, 156, 176, 180, 159, 55, 17]. usable [75]. use [116, 177, 214, 57, 113, 243, 216, 226, 123, 131, 161]. used [112]. User [126, 237, 71, 163, 337, 364, 301, 360, 338, 362, 326, 350, 152, 158, 162, 214, 57, 185, 181, 43, 88, 127, 21, 111, 266, 50, 68, 102, 156, 183, 161, 154, 65, 159, 47, 306, 358, 365]. User-Centered [163, 364, 57, 185, 21, 159]. user-contributed [154]. user-participants [162]. users [27, 69, 179, 106, 86, 178, 58]. Using [230, 251, 279, 259, 69, 200, 169]. Utah [237, 232, 234, 316, 241, 318]. UX [43, 50, 350, 345, 342, 364, 336, 160, 334, 184, 157, 147, 178, 333, 335, 340, 90, 353]. UXD [176].

Value [299, 120, 310, 126, 181, 4, 122]. value-proposition [126]. Verhulsdonck [364]. version [270]. via [248]. Vibrant [284]. video [178]. view [45]. viewers [182]. violence [253]. Violent [305]. virtues [241]. virtuous [200]. Visible [296, 147]. Vision [101]. Visual [344, 20, 8, 347, 95, 78, 56, 34]. visuality [204]. Visualization [309, 66, 132, 240, 89, 48]. Visualizer [309]. Visualizing [22, 223]. Visuals [347, 77]. Voice [293, 357]. Voices [337, 350, 149]. Void [337]. volunteer [99]. vs [353]. Vulnerability [344].

W [50]. WAC [306, 223]. Waltham [50]. Walton [256]. Warfare [264]. Warnick [257]. was [354]. Washington [72]. waters [86]. wearable [167, 165, 168, 169]. wearables [153]. web [185, 181, 31, 50, 289, 81, 82]. website [259]. websites [185, 181, 62]. WeChat [103]. Welchman [121]. well [103]. Wellbeing [288]. Wellness [288]. Wendy [305]. whistle [175]. whistle-blowing [175]. white [315]. Whittemore [108]. who [179]. Wicked [237]. Wild [237]. Wiley [233]. Wilfredo [291]. William [264]. within [236]. Women [149]. wonders [72]. WordPress [289]. Words [351]. Work [292, 328, 291, 60, 4, 128, 61, 267, 256, 247]. workflow [90]. working [69]. Workplace [281, 114, 351, 247]. Workshop [312]. World [295, 76, 107, 50]. writers [41]. Writing [35, 302, 269, 293, 364, 308, 360, 233, 310, 325, 146, 174, 16, 135, 133, 136, 96, 273, 308, 318]. wrong [55]. Wyatt [263].

XML [357, 129, 351].

Yak [117]. Yee [59]. Yee-Yin [59]. Yik [117]. Yin [59]. York [121, 42, 138, 49, 101, 58, 130, 222, 285]. you’re [12]. Yvonne [290].

Zealand [253]. Zines [361]. Zone [325].

References

Potts:2012:DDC

[1] Liza Potts and Michael Albers. Defining the design of communication. Communication Design Quarterly Review, 1 (1):3–7, September 2012. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Spinuzzi:2012:WCD

[2] Clay Spinuzzi. What is communication design? Communication Design Quarterly Review, 1(1):8–11, September 2012. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Swarts:2012:CD

[3] Jason Swarts. Communication design. Communication Design Quarterly Review, 1(1):12–15, September 2012. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hart-Davidson:2012:VCA

[4] William Hart-Davidson and Jeff Grabill. The value of computing, ambient data, ubiquitous connectivity for changing the work of communication designers. Communication Design Quarterly Review, 1 (1):16–22, September 2012. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hayhoe:2012:TFI

[5] George F. Hayhoe. Telling the future of information design. Communication Design Quarterly Review, 1(1): 23–26, September 2012. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

McNely:2012:BDS

[6] Brian McNely. Big data, situated people: humane approaches to communication design. Communication Design Quarterly Review, 1(1):27–30, September 2012. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Pierce:2012:DC

[7] Robert Pierce. Design of communication. Communication Design Quarterly Review, 1(1):31–36, September 2012. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Salvo:2012:VRB

[8] Michael J. Salvo. Visual rhetoric and big data: design of future communication. Communication Design Quarterly Review, 1(1):37–40, September 2012. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Potts:2013:NGD

[9] Liza Potts and Michael Albers. The next generation on design of communication. Communication Design Quarterly Review, 1(2):3–4, January 2013. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Koh:2013:CAL

[10] Jeffrey Tzu Kwan Valino Koh, Kening Zhu, Kasun Karunanayaka, Doros Polydorou, Roshan Lalintha Peiris, and Ryohei Nakatsu. Characterizing the analog-like and digital-like attributes of interactive systems. Communication Design Quarterly Review, 1(2): 8–36, January 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lindsley:2013:PID

[11] Tom Lindsley. Prefab interface development and the problem of ease. Communication Design Quarterly Review, 1 (2):37–49, January 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Harrison:2013:SYT

[12] Angela Harrison. I see you’re talking #HPV: communication patterns in the #HPV stream on Twitter. Communication Design Quarterly Review, 1(2): 50–51, January 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Potts:2013:SAE

[13] Liza Potts. SIGDOC at ATTW: editorial. Communication Design Quarterly Review, 1(3):3–4, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Pierce:2013:NC

[14] Rob Pierce. Notes from the chair. Communication Design Quarterly Review, 1(3):5–8, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Albers:2013:IPS

[15] Michael J. Albers. Introduction: Proceedings of Symposium on Communicating Complex Information. Communication Design Quarterly Review, 1 (3):9–11, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hailey:2013:RWP

[16] David E. Hailey, Jr. ReaderCentric writing for the prosumer marketplace: proposing a new, content-based information architecture model. Communication Design Quarterly Review, 1 (3):12–17, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Zobel:2013:ECU

[17] Gregory Zobel. Engaging complexity in usability through assemblage. Communication Design Quarterly Review, 1 (3):18–22, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Blythe:2013:DSM

[18] Stuart Blythe. Dynamic system models and the construction of complexity. Communication Design Quarterly Review, 1(3):23–27, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Shirey:2013:RCC

[19] Jenny Shirey, Ann Charng, and Quynh Nguyen. Researching and communicating the complexity of IT image management. Communication Design Quarterly Review, 1(3):28–33, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Meloncon:2013:VCE

[20] Lisa Meloncon. Visual communication in environmental health: methodological questions and compromises. Communication Design Quarterly Review, 1 (3):34–37, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Passera:2013:TCL

[21] Stefania Passera and Helena Haapio. Transforming contracts from legal rules to user-centered communication tools: a human-information interaction challenge. Communication Design Quarterly Review, 1(3):38–45, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Kain:2013:VCU

[22] Donna Kain and Michelle Covi. Visualizing complexity and uncertainty about climate change and sea level rise. Communication Design Quarterly Review, 1 (3):46–53, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Richardson:2013:AU

[23] Kevin H. Richardson. It’s not about usability. Communication Design Quarterly Review, 1(3):54–56, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Albers:2013:CCI

[24] Michael J. Albers. Cargo cults in information design. Communication Design Quarterly Review, 1(3):57–61, April 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Albers:2013:DCO

[25] Michael J. Albers. Design of communication open research questions: editorial. Communication Design Quarterly Review, 1(4):3–5, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Potts:2013:NC

[26] Liza Potts. Notes from the chair. Communication Design Quarterly Review, 1 (4):6–10, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Arduser:2013:PEU

[27] Lora Arduser. Produsers and end users: how social media impacts our students’ future research questions. Communication Design Quarterly Review, 1(4): 11–14, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Longo:2013:PRT

[28] Bernadette Longo, Nancy Coppola, Norbert Elliot, Andrew Klobucar, and Carol Johnson. A program of research for technical communication: adaptive learning. Communication Design Quarterly Review, 1(4):15–17, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Pflugfelder:2013:BDB

[29] Ehren Helmut Pflugfelder. Big data, big questions. Communication Design Quarterly Review, 1(4):18–21, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Jones:2013:RDB

[30] Dave Jones. From research to design: building knowledge so that we can build experiences. Communication Design Quarterly Review, 1(4):22–25, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Kalmbach:2013:IWN

[31] James Kalmbach. The invisible web and the need for new research methodologies. Communication Design Quarterly Review, 1(4):26–28, August 2013. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Davis:2013:ICP

[32] Marjorie T. Davis. Identifying core principles and expectations. Communication Design Quarterly Review, 1(4):29–30, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Walton:2013:NIC

[33] Rebecca Walton and Natasha N. Jones. Navigating increasingly crosscultural, cross-disciplinary, and crossorganizational contexts to support social justice. Communication Design Quarterly Review, 1(4):31–35, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Welhausen:2013:CMG

[34] Candice A. Welhausen. Chickens, MRIs, and graphics: creating visual information in scientific fields. Communication Design Quarterly Review, 1(4):36–39, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Abel:2013:WMT

[35] Scott Abel. Writing for machine translation. Communication Design Quarterly Review, 1(4):40–41, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Andersen:2013:ORQ

[36] Rebekka Andersen, Sid Benavente, Dave Clark, William Hart-Davidson, Carolyn Rude, and JoAnn Hackos. Open research questions for academics and industry professionals: results of a survey. Communication Design Quarterly Review, 1(4):42–49, August 2013. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Oswal:2013:EAP

[37] Sushil K. Oswal. Exploring accessibility as a potential area of research for technical communication: a modest proposal. Communication Design Quarterly Review, 1(4):50–60, August 2013. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Keller:2013:TDT

[38] Beth Keller. Tracing digital thyroid culture: poster. Communication Design Quarterly Review, 1(4):61, August 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Editors:2013:IDI

[39] Editors. Icon design to improve communication of health information to older adults. Communication Design Quarterly Review, 2(1):6–32, November 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Getto:2013:NKS

[40] Guiseppe Getto. Networked knowledges: student collaborative digital composing as communicative action. Communication Design Quarterly Review, 2(1): 33–58, November 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Aparicio:2013:TWL

[41] Manuela Aparicio. Technical writers @ Lisbon. Communication Design Quarterly Review, 2(1):59–60, November 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hennes:2013:BRB

[42] Jack Hennes. Book review: Rhetorical Accessability: At the Intersection of Technical Communication and Disability Studies, edited by Lisa Meloncon, Amityville, New York: Baywood, 2013. 247 pp. Communication Design Quarterly Review, 2(1):61–66, November 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Franklin:2013:BRB

[43] Nathan Franklin. Book review: The UX book: Process and guidelines for ensuring a quality user experience by Rex Hartson and Pardha A. Pyla, San Diego: Morgan Kaufmann. 2012. Communication Design Quarterly Review, 2(1): 67–72, November 2013. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Swarts:2014:MS

[44] Jason Swarts. The mobile situation. Communication Design Quarterly Review, 2(2):7–9, February 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Andersen:2014:TMI

[45] Rebekka Andersen. Toward a more integrated view of technical communication. Communication Design Quarterly Review, 2(2):10–16, February 2014. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lauer:2014:TCD

[46] Claire Lauer. Technology and communication design: crossroads and compromises. Communication Design Quarterly Review, 2(2):17–20, February 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Zhang:2014:BBH

[47] Tao Zhang, Ilana R. Barnes, and Marlen Promann. Building better help: user characteristics’ effect on library help design. Communication Design Quarterly Review, 2(2):21–27, February 2014. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Young:2014:RDV

[48] Justin Young and Charlie Potter. Remediation in data visualization: two examples of learning in real-time data processing environments. Communication Design Quarterly Review, 2(2): 29–37, February 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Keller:2014:BRB

[49] Beth Keller. Book review: Social media in disaster response: how experience architects can build for participation by L. Potts, (2013). New York, NY: Routledge. Communication Design Quarterly Review, 2(2):39–42, February 2014. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Ruszkiewicz:2014:BRB

[50] Sheryl Ruszkiewicz. Book review: Global UX: design and research in a connected world by W. Quesenbery and D. Szuc; Waltham, MA: Morgan Kaufmann and A web for everyone: designing accessible user experiences by S. Horton and W. Quesenbery; Brooklyn, NY: Rosenfeld media. Communication Design Quarterly Review, 2(2): 43–47, February 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Salvo:2014:WNE

[51] Michael Salvo. What’s in a name?: experience architecture rearticulates the humanities. Communication Design Quarterly Review, 2(3):6–9, May 2014. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Howard:2014:JMB

[52] Tharon Howard. Journey mapping: a brief overview. Communication Design Quarterly Review, 2(3):10–13, May 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Oswal:2014:PDB

[53] Sushil K. Oswal. Participatory design: barriers and possibilities. Communication Design Quarterly Review, 2(3):14– 19, May 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Shea:2014:CCT

[54] Marybeth Shea and Cameron Mozafari. Communicating complexity in transdisciplinary science teams for policy: applied stasis theory for organizing and assembling collaboration. Communication Design Quarterly Review, 2(3):20– 24, May 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Zhou:2014:UCW

[55] Quan Zhou. “That usability course”: what technical communication programs get wrong about usability and how to fix it. Communication Design Quarterly Review, 2(3):25–27, May 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Qian:2014:CAE

[56] Zhenyu Cheryl Qian, Yingjie Victor Chen, and Yinghuan Patty Peng. A comparative approach to enhance information interaction design of visual analytics systems. Communication Design Quarterly Review, 2(3):28–33, May 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Carlson:2014:LCS

[57] Clinton Carlson, Whitney Peake, and Jeff Joiner. Letting context speak: the use of co-creative, design-led, and usercentered design methods in the design of complex public communications. Communication Design Quarterly Review, 2 (3):34–39, May 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Labriola:2014:RCC

[58] Jack T. Labriola. Review of Crosscultural technology design: creating culture-sensitive technology for local users by Sun, H. (2012), New York, NY: Oxford University Press, Inc. Communication Design Quarterly Review, 2 (3):40–42, May 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2014:RCC

[59] Kirk St. Amant. Review of Crosscultural design for IT products and services by Pei-Luen Patrick Rau, Tom Plocher, & Yee-Yin Choong. (2013), CRC Press. Communication Design Quarterly Review, 2(3):43–45, May 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Andersen:2014:ECS

[60] Rebekka Andersen. The emergence of content strategy work and recommended resources. Communication Design Quarterly Review, 2(4):6–13, August 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

McNely:2014:KWK

[61] Brian J. McNely. Knowledge work, knowledge play: a heuristic approach to communication design for hybrid spaces. Communication Design Quarterly Review, 2(4):14–51, August 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Moeller:2014:PBN

[62] Marie Moeller. Pushing boundaries of normalcy: employing critical disability studies in analyzing medical advocacy websites. Communication Design Quarterly Review, 2(4):52–80, August 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Bethel:2014:RTG

[63] Chris Bethel. Review of “Topsight: a guide to studying, diagnosing, and fixing information flow in organizations” by Clay Spinuzzi, Amazon CreateSpace 2013, 978-1-4819-6006-9. Communication Design Quarterly Review, 2(4): 81–83, August 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Keller:2014:RPC

[64] Beth Keller. Review of “PowerPoint, Communication, and the Knowledge Society” by Hubert Knoblauch, Cambridge University Press 2013. Communication Design Quarterly Review, 2(4):84–86, August 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Wiley:2014:RUE

[65] Kristi Wiley. Review of “The user experience team of one: a research and design survival guide” by L. Buley Rosenfeld, Media 2013 978-1-933820-18-7. Communication Design Quarterly Review, 2 (4):87–89, August 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Aparicio:2014:DV

[66] Manuela Aparicio and Carlos J. Costa. Data visualization. Communication Design Quarterly Review, 3(1):7–11, November 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Daer:2014:RFH

[67] Alice R. Daer, Rebecca F. Hoffman, and Seth Goodman. Rhetorical functions of hashtag forms across social media applications. Communication Design Quarterly Review, 3(1):12–16, November 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Smith:2014:PAU

[68] Allegra W. Smith. Porn architecture: user tagging and filtering in two online pornography communities. Communication Design Quarterly Review, 3(1): 17–23, November 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Getto:2014:DGW

[69] Guiseppe Getto and Kirk St. Amant. Designing globally, working locally: using personas to develop online communication products for international users. Communication Design Quarterly Review, 3(1):24–46, November 2014. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Bethel:2014:BRD

[70] Chris Bethel. Book review: “The Digital Rights Movement: The Role of Technology in Subverting Digital Copyright” by Hector Postigo. The MIT Press, 2012. ISBN 978-0-262-01795-4. Communication Design Quarterly Review, 3(1): 47–48, November 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Bethel:2014:BRR

[71] Chris Bethel. Book review: “Responding to Technology — Resistance through Technology” (12–13), and “User Agency and Technology” (13–14). Communication Design Quarterly Review, 3(1): 49, November 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Butler:2014:BRM

[72] Janine M. Butler. Book review: “Morse, T. A. (2014). Signs and wonders: Religious rhetoric and the preservation of sign language”. Washington, DC: Gallaudet University Press. Communication Design Quarterly Review, 3(1): 50–53, November 2014. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hackos:2015:CTC

[73] JoAnn Hackos. Changing times: changing skills. Communication Design Quarterly Review, 3(2):7–12, February 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hailey:2015:WES

[74] David Hailey. To what extent should we re-examine our teaching? Communication Design Quarterly Review, 3(2): 13–19, February 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Howard:2015:PRU

[75] Tharon W. Howard. Are personas really usable? Communication Design Quarterly Review, 3(2):20–26, February 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Spyridakis:2015:INT

[76] Jan H. Spyridakis. Identifying new topics in TC curricula: preparing students for success in a changing world. Communication Design Quarterly Review, 3 (2):27–37, February 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2015:CCC

[77] Kirk St. Amant. Culture and the contextualization of care: a prototype-based approach to developing health and medical visuals for international audiences. Communication Design Quarterly Review, 3(2):38–47, February 2015. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

McNely:2015:TTS

[78] Brian McNely. Taking things seriously with visual research. Communication Design Quarterly Review, 3(2): 48–54, February 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

McDaniel:2015:UMA

[79] Rudy McDaniel. Understanding microinteractions as applied research opportunities for information designers. Communication Design Quarterly Review, 3(2):55–62, February 2015. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Oswal:2015:CUA

[80] Sushil K. Oswal. A conversation on usability and accessibility with Janice (Ginny) Redish. Communication Design Quarterly Review, 3(2):63–92, February 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hocutt:2015:RBI

[81] Daniel L. Hocutt. Review of Implementing Responsive Design: Building Sites for an Anywhere, Everywhere Web by Tim Kadlec, New Riders, 2013. ISBN 978-0-321-82168-3. Communication Design Quarterly Review, 3(2):

93–96, February 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Warren:2015:RBS

[82] Donovan Warren. Review of Mining the Social Web by Matthew A. Russell, Second edition. O’Reilly, 2013. ISBN 978-1-4493-6761-9. Communication Design Quarterly Review, 3(2):97–99, February 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Beale:2015:RPD

[83] Matthew Beale. Review of Playful Design: Creating Game Experiences in Everyday Interfaces. John Ferrara, Brooklyn, NY: Rosenfeld Media. 2012. ISBN: 978-1-933820-14-9. Communication Design Quarterly Review, 3(2): 100–103, February 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2015:AAC

[84] Kirk St. Amant. Aspects of access: considerations for creating health and medical content for international audiences. Communication Design Quarterly Review, 3(3):7–11, May 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Walton:2015:ERE

[85] Rebecca Walton and David Hailey. Evaluating the relevance of eBooks to corporate communication. Communication Design Quarterly Review, 3(3):12– 19, May 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Richards:2015:TWL

[86] Daniel Richards. Testing the waters: local users, sea level rise, and the productive usability of interactive geovisualizations. Communication Design Quarterly Review, 3(3):20–24, May 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Khataei:2015:PPB

[87] Amirsam Khataei and Ali Arya. Personalized presentation builder for persuasive communication. Communication Design Quarterly Review, 3(3):25– 32, May 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lanier:2015:PSU

[88] Clinton R. Lanier. Problem solving in user networks: complex communication issues and item-to-item collaborative filtering. Communication Design Quarterly Review, 3(3):33–39, May 2015. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Qian:2015:CEB

[89] Zhenyu Cheryl Qian and Yingjie Victor Chen. Communication and exchange between information visualization and industrial design. Communication Design Quarterly Review, 3(3):40–48, May 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Wiley:2015:UWB

[90] Kristi Wiley and Guiseppe Getto. A UX workflow for building awesome application. Communication Design Quarterly Review, 3(3):49–52, May 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Zhou:2015:SFE

[91] Quan Zhou. Strategy first, execution second: teaching design strategy in technical communication. Communication Design Quarterly Review, 3(3):53–55, May 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Fanfarelli:2015:UDB

[92] Joseph Fanfarelli, Stephanie Vie, and Rudy McDaniel. Understanding digital badges through feedback, reward, and narrative: a multidisciplinary approach to building better badges in social environments. Communication Design Quarterly Review, 3(3):56–60, May 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Meloncon:2015:SII

[93] Lisa Meloncon and Erin A. Frost. Special issue introduction: Charting an emerging field: the rhetorics of health and medicine and its importance in communication design. Communication Design Quarterly Review, 3(4):7–14, August 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Connellan:2015:LLH

[94] Kathleen Connellan, Damien W. Riggs, and Clemence Due. Light lies: how glass speaks. Communication Design Quarterly Review, 3(4):15–24, August 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lazard:2015:HFI

[95] Allison J. Lazard and Michael S. Mackert. e-health first impressions and visual evaluations: key design principles for attention and appeal. Communication Design Quarterly Review, 3(4):25–34, August 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Mogull:2015:PCW

[96] Scott A. Mogull and Deborah Balzhiser. Pharmaceutical companies are writing the script for health consumerism. Communication Design Quarterly Review, 3 (4):35–49, August 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Burleson:2015:HMH

[97] Debra Burleson. The hospitalist model: are hospitals informing patients? Communication Design Quarterly Review, 3 (4):50–60, August 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Novotny:2015:RGR

[98] Maria Novotny. reVITALize gynecology: reimagining apparent feminism’s methodology in participatory health intervention projects. Communication Design Quarterly Review, 3(4):61–74, August 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Kuehl:2015:DPC

[99] Rebecca A. Kuehl and Jenn Anderson. Designing public communication about doulas: analyzing presence and absence in promoting a volunteer doula program. Communication Design Quarterly Review, 3(4):75–84, August 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Avtgis:2015:AAT

[100] Theodore A. Avtgis, David Kappel, E. Phillips Polack, Alison Wilson, and Jennifer Knight. Assessing the accuracy of trauma patient prioritization: communication design of the M.I.S.E.R information system protocol and communication channel during crisis communication exchanges. Communication Design Quarterly Review, 3(4):85–90, August 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Kessler:2015:RRF

[101] Molly Kessler. Review of “Rhetoric in the Flesh: Trained Vision, Technical Expertise, and the Gross Anatomy Lab. by T. Kenny Fountain” New York, NY: Routledge, 2014. Communication Design Quarterly Review, 3(4):91–96, August 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2015:ISI

[102] Kirk St. Amant. Introduction to the special issue: Cultural considerations for communication design: integrating ideas of culture, communication, and context into user experience design. Communication Design Quarterly Review, 4(1): 6–22, November 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Wang:2015:CDW

[103] Xiaobo Wang and Baotong Gu. The communication design of WeChat: ideological as well as technical aspects of social media. Communication Design Quarterly Review, 4(1):23–35, November 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Dutta:2015:DDM

[104] Uttaran Dutta and Swayang Das. The digital divide at the margins: codesigning information solutions to address the needs of indigenous populations of rural India. Communication Design Quarterly Review, 4(1):36– 48, November 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

McDaniel:2015:CCC

[105] Rudy McDaniel and Lanlan Kuang.

Cross-cultural cinematic communication: learning from the information design process for a Sino–American film competition. Communication Design Quarterly Review, 4(1):49–60, November 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Peter:2015:DHD

[106] Hilary Sarat-St. Peter. Designing with HDR data: what the human development report can tell us about international users. Communication Design Quarterly Review, 4(1):60–72, November 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Getto:2015:RWG

[107] Guiseppe Getto. Review of “World 3.0:

Global prosperity and how to achieve it by P. Ghemawat”, Harvard Business Review Press 2011. Communication Design Quarterly Review, 4(1):73–76, November 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lauren:2015:RRM

[108] Benjamin Lauren. Review of “Rhetorical memory: a study of technical communication and information management by S. Whittemore”, University of Chicago Press 2015. Communication Design Quarterly Review, 4(1):77–80, November 2015. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Amant:2016:ONS

[109] Kirk St. Amant. Online networks, social media, and communication design. Communication Design Quarterly Review, 4(2):10–11, March 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Frith:2016:FDC

[110] Jordan Frith. Forum design and the changing landscape of crowd-sourced help information. Communication Design Quarterly Review, 4(2):12–22, March 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Petersen:2016:EUD

[111] Emily January Petersen. Empathetic user design: understanding and living the reality of an audience. Communication Design Quarterly Review, 4 (2):23–36, March 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lam:2016:SHD

[112] Chris Lam and Mark A. Hannah. The social help desk: examining how Twitter is used as a technical support tool. Communication Design Quarterly Review, 4 (2):37–51, March 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Faris:2016:ESS

[113] Michael J. Faris and Kristen R. Moore. Emerging scholars and social media use: a pilot study of risk. Communication Design Quarterly Review, 4(2):52–63, March 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Gunning:2016:RAE

[114] Sarah K. Gunning. Review of “All edge: Inside the new workplace networks”, by Spinuzzi, C., University of Chicago Press: Chicago (2015). Communication Design Quarterly Review, 4 (2):64–68, March 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Pigg:2016:RII

[115] Stacey L. Pigg. Review of “Intertwingled: Information changes everything”, by Morville, P., Semantic Studios, Ann Arbor, MI (2014). Communication Design Quarterly Review, 4 (2):69–72, March 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Amant:2016:AAC

[116] Kirk St. Amant. Aspects of awareness:

considerations for social media use in the modern context. Communication Design Quarterly Review, 4(2b):8–10, June 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

West:2016:YYK

[117] Sara West. Yik Yak and the knowledge community. Communication Design Quarterly Review, 4(2b):11–21, June 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hinson:2016:FIT

[118] Katrina L. Hinson. Framing illness through Facebook enabled online support groups. Communication Design Quarterly Review, 4(2b):22–31, June 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hopton:2016:SSS

[119] Sarah-Beth Hopton and R. Mitchell Parry. Saving the sea, socially: measuring the relationship between content and gesture on Facebook. Communication Design Quarterly Review, 4(2b):32– 43, June 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Jones:2016:RME

[120] Dave Jones. Review of “Mapping Experiences: A Guide to Creating Value through Journeys, Blueprints, and Diagrams, by Kalbach, J.,” Sebastopol, CA: O’Reilly Media (2016). Communication Design Quarterly Review, 4(2b):44–48, June 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Andrews:2016:RMC

[121] Christopher Andrews. Review of “Managing Chaos: Digital Governance by Design”, by Welchman, L., New York: Rosenfeld Media (2015). Communication Design Quarterly Review, 4(2b):49– 53, June 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2016:RCN

[122] Kirk St. Amant. Re-considering the nature of value in communication design. Communication Design Quarterly Review, 4(3):4–8, September 2016. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Oberg:2016:ECT

[123] Lena-Maria Öberg. Examining the context of technical information use: special section introduction. Communication Design Quarterly Review, 4(3): 9–11, September 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Soderlund:2016:WIS

[124] C. Söderlund and J. Lundin. What is an information source?: information design based on information source selection behavior. Communication Design Quarterly Review, 4(3):12–19, September 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lowgren:2016:TCP

[125] Jonas Löwgren. Technical communication practices in the collaborative mediascape: a case study in media structure transformation. Communication Design Quarterly Review, 4(3):20–25, September 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Acharya:2016:UVU

[126] Keshab R. Acharya. User value and usability in technical communication: a value-proposition design model. Communication Design Quarterly Review, 4 (3):26–34, September 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Opel:2016:DOR

[127] Dawn Opel. Designing online resources for safety net healthcare providers: users’ needs and the evidence-based medicine paradigm. Communication Design Quarterly Review, 4(3):35–45, September 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lauer:2016:ERC

[128] Claire Lauer. Editorial re-considering research: why we need to adopt a mixedmethods approach to our work. Communication Design Quarterly Review, 4(3): 46–50, September 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Moxley-Kelly:2016:RLT

[129] Sean Moxley-Kelly. Review of “The Language of Technical Communication,” by Gallon, R. (2016). Laguna Hills, CA: XML Press. Communication Design Quarterly Review, 4(3):51–55, September 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Strantz:2016:RMS

[130] Adam Strantz. Review of “The Mobile Story: Narrative Practices with Locative Technologies,” by Farman, J. (Ed.). (2014). New York, NY: Routledge. Communication Design Quarterly Review, 4 (3):56–61, September 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2016:FCU

[131] Kirk St. Amant. Of form, context, and use. Communication Design Quarterly Review, 4(4):4–6, December 2016. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hepworth:2016:BDV

[132] Katherine Hepworth. Big data visualization: promises & pitfalls. Communication Design Quarterly Review, 4(4): 7–19, December 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hewett:2016:RWD

[133] Beth L. Hewett. Reading, writing, and digital composition: reintegrating constituent literacies in online settings. Communication Design Quarterly Review, 4(4):20–35, December 2016. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

deWinter:2016:MCM

[134] Jennifer deWinter, Carly A. Kocurek, and Stephanie Vie. Managing community managers: social labor, feminized skills, and professionalization. Communication Design Quarterly Review, 4(4): 36–45, December 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Harris:2016:UTD

[135] Heidi Skurat Harris and Michael Greer.

Over, under, or through: design strategies to supplement the LMS and enhance interaction in online writing courses. Communication Design Quarterly Review, 4(4):46–54, December 2016. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Licastro:2016:PMW

[136] Amanda M. Licastro. The problem of multimodality: what data-driven research can tell us about online writing practices. Communication Design Quarterly Review, 4(4):55–73, December 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Kiwanuka-Tondo:2016:LCS

[137] James Kiwanuka-Tondo and Keon Mandell Pettiway. Localizing complex scientific communication: a SWOT analysis and multi-sectoral approach of communicating climate change. Communication Design Quarterly Review, 4(4): 74–85, December 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hocutt:2016:RCM

[138] Daniel L. Hocutt. Review of “Communicating Mobility and Technology: A Material Rhetoric for Persuasive Transportation,” by Pflugfelder, E. H. (2017). New York: Routledge, 2017. Communication Design Quarterly Review, 4(4): 86–92, December 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Shivers-McNair:2016:RRC

[139] Ann Shivers-McNair. Review of “Risk Communication and Miscommunication: Case Studies in Science, Technology, Engineering, Government, and Community Organizations,” by Boiarsky, C. (2016). Boulder, CO: University of Colorado Press. Communication Design Quarterly Review, 4(4): 93–98, December 2016. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2017:CCCa

[140] Kirk St. Amant. Of content, context, and conveyance: editorial. Communication Design Quarterly Review, 5(1):4–7, April 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Graham:2017:DLT

[141] S. Scott Graham. Data and lore in technical communication research: guest editorial. Communication Design Quarterly Review, 5(1):8–25, April 2017. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Zaad:2017:AIC

[142] Lambert Zaad, Dick Lenior, Thea van der Geest, and Els van der Pool. Analyzing information in complex collaborative tasks. Communication Design Quarterly Review, 5(1):26–42, April 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Li:2017:CID

[143] Mingran Li, Ruimin Gao, Xinghe Hu, and Yingjie Chen. Comparing Info-Vis designs with different information architecture for communicating complex information. Communication Design Quarterly Review, 5(1):43–56, April 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Flanagan:2017:HCT

[144] Suzan Flanagan and Guiseppe Getto.

Helping content: a three-part approach to content strategy with nonprofits. Communication Design Quarterly Review, 5(1):57–70, April 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

King:2017:DSC

[145] Abigail Selzer King, Kristen R. Moore, Ashley Hardage Edlin, and Sophie Frankel. Drawing strategies for communication planning: a rationale and exemplar of the geometric page form (GPF) approach. Communication Design Quarterly Review, 5(1):71–79, April 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Bourelle:2017:DOW

[146] Tiffany Bourelle, Angela Clark-Oates, and Andrew Bourelle. Designing online writing classes to promote multimodal literacies: five practices for course design. Communication Design Quarterly Review, 5(1):80–88, April 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Rose:2017:MPL

[147] Emma Rose and Josh Tenenberg. Making practice-level struggles visible: researching UX practice to inform pedagogy. Communication Design Quarterly Review, 5(1):89–97, April 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Deaton:2017:RSC

[148] Phillip J. Deaton. Reading sounds: closed-captioned media and popular culture: book review. Communication Design Quarterly Review, 5(1):98–101, April 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Keller:2017:WVM

[149] Beth Keller. Women’s voices in management: identifying innovative and responsible solutions: book review. Communication Design Quarterly Review, 5 (1):102–105, April 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Heifferon:2017:NTP

[150] Barbara Heifferon. New technologies, patient experience, theoretical approaches and heuristics in RHM: guest editorial. Communication Design Quarterly Review, 5(2):4–18, July 2017. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Meloncon:2017:PED

[151] Lisa K. Meloncon. Patient experience design: expanding usability methodologies for healthcare. Communication Design Quarterly Review, 5(2):19–28, July 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Angeli:2017:RPH

[152] Elizabeth L. Angeli and Christina D.

Norwood. Responding to public health crises: bridging collective mindfulness and user experience to create communication interventions. Communication Design Quarterly Review, 5(2):29–39, July 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Jones:2017:DPH

[153] John Jones, Catherine Gouge, and Mariah Crilley. Design principles for health wearables. Communication Design Quarterly Review, 5(2):40–50, July 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Welhausen:2017:YOR

[154] Candice A. Welhausen. At your own risk: user-contributed flu maps, participatory surveillance, and an emergent DIY risk assessment ethic. Communication Design Quarterly Review, 5(2):51– 61, July 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2017:CCCb

[155] Kirk St. Amant. The cultural context of care in international communication design: a heuristic for addressing usability in international health and medical communication. Communication Design Quarterly Review, 5(2):62–70, July 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2017:EIR

[156] Kirk St. Amant. Editor’s introduction: Reflecting on and re-thinking usability and user experience design. Communication Design Quarterly Review, 5(3): 4–9, November 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Robinson:2017:PPF

[157] Joy Robinson, Candice Lanius, and Ryan Weber. The past, present, and future of UX empirical research. Communication Design Quarterly Review, 5(3): 10–23, November 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Applen:2017:DHM

[158] J. D. Applen and Sonia H. Stephens. Digital humanities, middleware, and user experience design for public health applications. Communication Design Quarterly Review, 5(3):24–34, November 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Williamson:2017:CRA

[159] Bill Williamson and Scott J. Kowalewski. Cultivating a rhetoric of advocacy for usability studies and user-centered design. Communication Design Quarterly Review, 5(3):35–47, November 2017. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Mara:2017:FNU

[160] Andrew Mara. Framework negotiation and UX design. Communication Design Quarterly Review, 5(3):48–54, November 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Verhulsdonck:2017:DGM

[161] Gustav Verhulsdonck. Designing for global mobile: considering user experience mapping with infrastructure, global openness, local user contexts and local cultural beliefs of technology use. Communication Design Quarterly Review, 5 (3):55–62, November 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Bartolotta:2017:EOD

[162] Joseph Bartolotta, Julianne Newmark, and Tiffany Bourelle. Engaging with online design: undergraduate userparticipants and the practice-level struggles of usability learning. Communication Design Quarterly Review, 5(3): 63–72, November 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Cosgrove:2017:RFU

[163] Samantha Cosgrove. Review of “Fundamentals of User-Centered Design: A Practical Approach,” by Still, B., & Crane, K. (2017). Boca Raton, FL: CRC Press. Communication Design Quarterly Review, 5(3):74–77, November 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Rowan:2017:REP

[164] Robert M. Rowan. Review of “Exploratory Programming for the Arts and Humanities,” by Montfort, N. (2016). Cambridge, MA: MIT Press. Communication Design Quarterly Review, 5(3): 78–82, November 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Jones:2017:GEI

[165] John Jones and Catherine C. Gouge. Guest editors’ introduction wearable technologies and communication design. Communication Design Quarterly Review, 5(4):4–14, December 2017. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Amidon:2017:SGD

[166] Timothy R. Amidon, Elizabeth A. Williams, Tiffany Lipsey, Randy Callahan, Gary Nuckols, and Spencer Rice. Sensors and gizmos and data, oh my: informating firefighters’ personal protective equipment. Communication Design Quarterly Review, 5(4):15–30, December 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Arduser:2017:IPD

[167] Lora Arduser. Impatient patients: a DIY usability approach in diabetes wearable technologies. Communication Design Quarterly Review, 5(4):31– 39, December 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Kennedy:2017:DHM

[168] Krista Kennedy. Designing for humanmachine collaboration: smart hearing aids as wearable technologies. Communication Design Quarterly Review, 5(4): 40–51, December 2017. CODEN ????

ISSN 2166-1200 (print), 2166-1642 (electronic).

Moore:2017:CSH

[169] Kristen R. Moore, Natasha Jones, Bailey S. Cundiff, and Leah Heilig. Contested sites of health risks: using wearable technologies to intervene in racial oppression. Communication Design Quarterly Review, 5(4):52–60, December 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Welhausen:2017:QMF

[170] Candice A. Welhausen. Quantifiable me:

fitness and health trackers and the trope of holisticism. Communication Design Quarterly Review, 5(4):61–71, December 2017. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Richards:2018:PFP

[171] Daniel Richards. Proceedings from and future plans for the Symposium for Communicating Complex Information (SCCI): Guest Editor’s introduction. Communication Design Quarterly Review, 6(1):4–8, March 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Youngblood:2018:SIA

[172] Susan A. Youngblood. Site identity, artifact duplication, and disambiguation in Alabama Local Emergency Management Agencies (LEMAs). Communication Design Quarterly Review, 6 (1):9–15, March 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Opel:2018:WOH

[173] Dawn S. Opel. What is “Obamacare”?: health literacy, e-commerce, and the Affordable Care Act’s online content. Communication Design Quarterly Review, 6(1):16–25, March 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

DeTora:2018:PTC

[174] Lisa DeTora. Principles of technical communication and design can enrich writing practice in regulated contexts. Communication Design Quarterly Review, 6(1):26–34, March 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Ding:2018:CCW

[175] Huiling Ding. Cross-cultural whistleblowing in an emerging outbreak: revealing health risks through tactic communication and rhetorical hijacking. Communication Design Quarterly Review, 6(1):35–44, March 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

StAmant:2018:RRU

[176] Kirk St. Amant. Reflexes, reactions, and usability: examining how prototypes of place can enhance UXD practices. Communication Design Quarterly Review, 6 (1):45–53, March 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Amant:2018:MCC

[177] Kirk St. Amant. Mapping the complex context(s) of use: editorial. Communication Design Quarterly Review, 6(2):4–8, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Rose:2018:PVM

[178] Emma Rose and Alison Cardinal. Participatory video methods in UX: sharing power with users to gain insights into everyday life. Communication Design Quarterly Review, 6(2):9–20, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hutter:2018:PIA

[179] Liz Hutter and Halcyon M. Lawrence.

Promoting inclusive and accessible design in usability testing: a teaching case with users who are deaf. Communication Design Quarterly Review, 6(2): 21–30, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Walkup:2018:CYP

[180] Katie Lynn Walkup. Connect with your patients, not the screen: usability claims in electronic health records. Communication Design Quarterly Review, 6(2): 31–40, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Everett:2018:GEG

[181] Heidi L. Everett. Is good enough good enough?: negotiating web user value judgments of small businesses based on poorly designed websites. Communication Design Quarterly Review, 6(2): 41–56, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Richards:2018:CLP

[182] Daniel P. Richards. Not a cape, but a life preserver: the importance of designer localization in interactive sea level rise viewers. Communication Design Quarterly Review, 6(2):57–69, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Thominet:2018:HOU

[183] Luke Thominet. How to be open: user experience and technical communication in an emerging game development methodology. Communication Design Quarterly Review, 6(2):70–82, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Reimer:2018:CCC

[184] Cody Reimer. Contextual cropping, collateral data: screenshot methods for UX research. Communication Design Quarterly Review, 6(2):83–92, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Cosgrove:2018:EUU

[185] Samantha Cosgrove. Exploring usability and user-centered design through emergency management websites: advocating responsive web design. Communication Design Quarterly Review, 6(2): 93–102, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Clinkenbeard:2018:MCA

[186] Mary Clinkenbeard. Multimodal conversation analysis and usability studies: exploring human-technology interactions in multiparty contexts. Communication Design Quarterly Review, 6(2): 103–113, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Rubens:2018:BR

[187] Amy Rubens. Book review. Communication Design Quarterly Review, 6(2): 114–118, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Walwema:2018:BR

[188] Josephine Walwema. Book review. Communication Design Quarterly Review, 6 (2):119–123, Summer 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Andersen:2018:EPP

[189] Rebekka Andersen and Carlos Evia. Editorial: perspectives on preparing technical communication professionals for today and the future. Communication Design Quarterly Review, 6(3):4– 13, September 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Shalamova:2018:ESS

[190] Nadya Shalamova, Tammy Rice-Bailey, and Katherine Wikoff. Evolving skill sets and job pathways of technical communicators. Communication Design Quarterly Review, 6(3):14–24, September 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Carnegie:2018:RCC

[191] Teena A. M. Carnegie and Kate Crane.

Responsive curriculum change: going beyond occupation demands. Communication Design Quarterly Review, 6(3): 25–31, September 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Swope:2018:IAW

[192] Amber Swope. Information architects: what they do and how to become one. Communication Design Quarterly Review, 6(3):32–43, September 2018. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Duin:2018:CCL

[193] Ann Hill Duin and Jason Chew Kit Tham. Cultivating code literacy: course redesign through advisory board engagement. Communication Design Quarterly Review, 6(3):44–58, September 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lauren:2018:PCD

[194] Benjamin Lauren. Preparing communication design students as facilitators: a primer for rethinking coursework in project management. Communication Design Quarterly Review, 6(3): 59–65, September 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Atkins:2018:PSI

[195] Anthony T. Atkins and Colleen A. Reilly. Pedagogical strategies for integrating SEO into technical communication curricula. Communication Design Quarterly Review, 6(3):66–73, September 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Mallette:2018:TPN

[196] Jennifer C. Mallette and Megan Gehrke. Theory to practice: negotiating expertise for new technical communicators. Communication Design Quarterly Review, 6(3):74–83, September 2018. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Martin:2018:BR

[197] Dan Martin. Book review. Communication Design Quarterly Review, 6(3): 84–88, September 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Wisniewski:2018:BR

[198] Elaine Wisniewski. Book review. Communication Design Quarterly Review, 6 (3):89–93, September 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Zdenek:2018:GEI

[199] Sean Zdenek. Guest Editor’s introduction: reimagining disability and accessibility in technical and professional communication. Communication Design Quarterly Review, 6(4):4–11, December 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Huntsman:2018:CVC

[200] Sherena Huntsman, Jared S. Colton, and Christopher Phillips. Cultivating virtuous course designers: using technical communication to reimagine accessibility in higher education. Communication Design Quarterly Review, 6(4): 12–23, December 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Garrison:2018:TLR

[201] Kevin Garrison. Theorizing lip reading as interface design: the gadfly of the gaps. Communication Design Quarterly Review, 6(4):24–34, December 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Gonzales:2018:DII

[202] Laura Gonzales. Designing for intersectional, interdependent accessibility: a case study of multilingual technical content creation. Communication Design Quarterly Review, 6(4):35–45, December 2018. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Hopton:2019:GEI

[203] Sarah Beth Hopton. Guest editor’s introduction: the revenge of Plato’s pigs. Communication Design Quarterly Review, 7(1):4–8, April 2019. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Eichberger:2019:MSS

[204] Ryan Eichberger. Maps, silence, and Standing Rock: seeking a visuality for the age of environmental crisis. Communication Design Quarterly Review, 7 (1):9–21, April 2019. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Cagle:2019:SDD

[205] Lauren E. Cagle and Carl Herndl. Shades of denialism: discovering possibilities for a more nuanced deliberation about climate change in online discussion forums. Communication Design Quarterly Review, 7(1):22–39, April 2019. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

George:2019:CAR

[206] Barbara George. Communicating activist roles and tools in complex energy deliberation. Communication Design Quarterly Review, 7(1):40–53, April 2019. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Ballentine:2019:RRH

[207] Brian Ballentine. Rhetoric, risk, and hydraulic fracturing: one landowner’s perspective. Communication Design Quarterly Review, 7(1):54–63, April 2019. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Vernon:2019:CPB

[208] Laura Vernon. Crossing political borders: how a grassroots environmental group influenced a change in public policy. Communication Design Quarterly Review, 7(1):64–72, April 2019. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Pflugfelder:2019:RSN

[209] Ehren Helmut Pflugfelder. Risk selfies and nonrational environmental communication. Communication Design Quarterly Review, 7(1):73–84, April 2019. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Dong:2019:EDC

[210] Lin Dong. Earth discourses: constructing risks and responsibilities in Chinese state and social media. Communication Design Quarterly Review, 7(1):85–99, April 2019. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Ross:2019:EI

[211] Derek G. Ross. Editor’s introduction.

Communication Design Quarterly Review, 7(2):4–5, July 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Albers:2019:E

[212] Michael Albers. Editorial. Communication Design Quarterly Review, 7(2):6, July 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Rogers:2019:TDB

[213] Ryan Rogers and Laura Dunlow. Testing the difference between appearance and ability customization. Communication Design Quarterly Review, 7(2):7– 16, July 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Bivens:2019:RHD

[214] Kristin Marie Bivens. Reducing harm by designing discourse and digital tools for opioid users’ contexts: the Chicago Recovery Alliance’s community-based context of use and PwrdBy’s technologybased context of use. Communication Design Quarterly Review, 7(2):17–27, July 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Hierro:2019:DPC

[215] Victor Del Hierro. DJs, playlists, and community: imagining communication design through hip hop. Communication Design Quarterly Review, 7(2):28– 39, July 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Meng:2019:HDU

[216] Michael Meng, Stephanie Steinhardt, and Andreas Schubert. How developers use API documentation: an observation study. Communication Design Quarterly Review, 7(2):40–49, July 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Edenfield:2019:QCD

[217] Avery C. Edenfield. Queering consent:

design and sexual consent messaging. Communication Design Quarterly Review, 7(2):50–63, July 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Potts:2019:RRE

[218] L. Potts, M. J. Salvo, and Leslie Hankey. Review of “Rhetoric and experience architecture”, Parlor Press: Liza Potts and Michael Salvo. Communication Design Quarterly Review, 7(2):64–65, July 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Edenfield:2019:BRE

[219] Avery Edenfield. From the book review editor. Communication Design Quarterly Review, 7(3):4, September 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Colby:2019:GDD

[220] Richard Colby and Rebekah Shultz Colby. Game design documentation: four perspectives from independent game studios. Communication Design Quarterly Review, 7(3):5– 15, September 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Bartolotta:2019:UTO

[221] Joseph Bartolotta. Usability testing for oppression. Communication Design Quarterly Review, 7(3):16–29, September 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Zobel:2019:RAO

[222] Gregory Zobel. Review of “Algorithms of oppression: how search engines reinforce racism,” by Noble, S. U. (2018). New York, New York: NYU Press. Communication Design Quarterly Review, 7 (3):30–31, September 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Zobel:2019:RNS

[223] Gregory Zobel. Review of “Network sense: methods for visualizing a discipline,” by Mueller, D. N. (2017). Fort Collins, Colorado: WAC Clearinghouse. Communication Design Quarterly Review, 7(3):32–33, September 2019. ISSN 2166-1200 (print), 2166-1642 (electronic).

Richards:2020:USC

[224] Daniel P. Richards and Derek G. Ross. Updates from SIGDOC and CDQ: editorial. Communication Design Quarterly Review, 8(1):4, May 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3375134.3375139.

Stephens:2020:SMS

[225] Sonia H. Stephens and Daniel P. Richards. Story mapping and sea level rise: listening to global risks at street level. Communication Design Quarterly Review, 8(1):5–18, May 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3375134.
3375135.

Miller:2020:SMP

[226] Jennifer Roth Miller, Brandy Dieterle, Jennifer deWinter, and Stephanie Vie. Social media in professional, technical, and scientific communication programs: a heuristic to guide future use. Communication Design Quarterly Review, 8 (1):19–34, May 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3375134.3375136.

Cheek:2020:BRD

[227] Ryan Cheek. Book review of “Design, ecology, politics: towards the ecocene” by Joanna Boehnert (2018). Bloomsbury Academic. Communication Design Quarterly Review, 8(1): 35–36, May 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3375134.3375137.

Browning:2020:BRB

[228] Ella R. Browning. Book review of “Bodies in flux: scientific methods for negotiating medical uncertainty” by Christa Teston (2017). University of Chicago Press. Communication Design Quarterly Review, 8(1): 37–39, May 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3375134.3375138.

Richards:2020:OSS

[229] Dan Richards, Sarah Read, Susan Youngblood, Emma Rose, and Derek G. Ross. Official statement from SIG-DOC: a response to injustice. Communication Design Quarterly Review, 8 (2):4–5, August 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3394264.3394267.

Applen:2020:UBI

[230] J. D. Applen. Using Bayesian induction methods in risk assessment and communication. Communication Design Quarterly Review, 8(2):

6–15, August 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3394264.3394265.

Oppegaard:2020:PPA

[231] Brett Oppegaard. Prototyping and public art: design and field studies in locative media. Communication Design Quarterly Review, 8(2): 16–27, August 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3394264.3394266.

Clark:2020:RTP

[232] Tracy Clark. Review of “Teaching Professional and Technical Communication: A Practicum in a Book” by Tracy Bridgeford, Bridgeford, T. (2018). Teaching professional and technical communication: a practicum in a book. Utah State University Press. Communication Design Quarterly Review, 8 (2):28–29, August 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3394264.3394268.

McPherson:2020:RIG

[233] Cynthia McPherson. Review of “The IEEE Guide to Writing in the Engineering and Technical Fields” by David Kmiec and Bernadette Longo, Kmiec, D. & Longo, B. (2017). The IEEE guide to writing in the engineering and technical fields. John Wiley & Sons, Inc. Communication Design Quarterly Review, 8 (2):30–31, August 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3394264.3394269.

Mitchell:2020:RKT

[234] Claudia Mitchell. Review of “Key Theoretical Frameworks: Teaching Technical Communication in the Twenty-First Century” by Angela M. Haas and Michelle F. Eble, Haas, A. M., & Eble, M. F. (2018). Key theoretical frameworks: Teaching technical communication in the twenty-first century. Utah State University. Communication Design Quarterly Review, 8(2): 32–33, August 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/abs/10.1145/3394264.3394270.

Lauer:2020:ITD

[235] Claire Lauer. Implementing a transactional design model to ensure the mindful development of public-facing science communication projects. Communication Design Quarterly Review, 8(3):4– 15, November 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3410430.3436988.

Slotkin:2020:ACP

[236] Alexander Slotkin. Along the cow path: technical communication within a Jewish Cemetery. Communication Design Quarterly Review, 8(3):16– 25, November 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3410430.3436989.

Berger:2020:RWI

[237] Arthur Berger. Review of “Wicked, Incomplete, and Uncertain: User Support in the Wild and the Role of Technical Communication by Jason Swarts (2018),” Utah State University Press.

Communication Design Quarterly Review, 8(3):26–27, November 2020. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/3410430.
                                        3436990.

Pfannenstiel:2020:RCS

[238] A. Nicole Pfannenstiel. Review of “Content Strategy in Technical Communication by Guiseppe Getto, Jack T. Labriola, and Sheryl Ruszkiewicz (Eds.). (2020),” Content strategy in technical communication. Routledge. Communication Design Quarterly Review, 8(3): 28–29, November 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3410430.3436991.

Cheek:2020:PTC

[239] Ryan Cheek. Political technical communication and ideographic communication design in a pre-digital congressional campaign. Communication Design Quarterly Review, 8(4):4– 14, December 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3431932.3431933.

Olman:2020:HCH

[240] Lynda Olman and Danielle DeVasto. Hybrid collectivity: hacking environmental risk visualization for the Anthropocene. Communication Design Quarterly Review, 8(4):15–28, December 2020. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3431932.3431934.

Vance:2020:RRT

[241] Bremen Vance and Lauren Malone. Review of “Rhetoric technology and the virtues by Jared S. Colton and Steve Holmes,”[ Colton, J. S., & Holmes, S. (2018). Rhetoric, technology, and the virtues. Utah State University Press]. Communication Design Quarterly Review, 8(4):29–30, December 2020. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3431932.
3431935.

Butts:2021:DME

[242] Shannon Butts and Madison Jones.

Deep mapping for environmental communication design. Communication Design Quarterly Review, 9(1):4–19, March 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3437000.3437001.

Itchuaqiyaq:2021:DDC

[243] Cana Uluak Itchuaqiyaq and Breeanne Matheson. Decolonizing decoloniality: considering the (mis)use of decolonial frameworks in TPC scholarship. Communication Design Quarterly Review, 9 (1):20–31, March 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3437000.3437002.

Ibrahim:2021:RNT

[244] Mai Ibrahim. Review of “Nihilism and technology by Nolen Gertz,” [Gertz, N. (2018). Nihilism and technology. Rowman & Littlefield]. Communication Design Quarterly Review, 9(1): 32–34, March 2021. CODEN ????

ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3437000.3437003.

West:2021:RRH

[245] Temple West. Review of “Rhetoric of health and medicine as/is: Theories and approaches for the field by Lisa Melonçon, S. Scott Graham, Jenell Johnson, John A. Lynch, and Cynthia Ryan,” [Melonçon, L. Graham, S.S, Johnson, J., Lynch, J., & Ryan, S. (Eds). (2020). Rhetoric of health and medicine as/is: Theories and approaches for the field. The Ohio State University Press. https://doi.org/10.26818/9780814214466]. Communication Design Quarterly Review, 9(1):35–36, March 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

org/doi/10.1145/3437000.3437004.

Cogbill-Seiders:2021:RSC

[246] Elisa Cogbill-Seiders. Review of “The science of communicating science by Craig Cormick,” [Cormick, C. (2019). The science of communicating science. CSIRO publishing]. Communication Design Quarterly Review, 9(1): 37–38, March 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3437000.3437005.

Weems:2021:RRW

[247] Elizabeth E. Weems. Review of “Rhetorical work in emergency medical services: Communicating in the unpredictable workplace by Elizabeth Angeli,” [Angeli, E. L. (2019). Rhetorical work in emergency medical services: communicating in the unpredictable workplace. Routledge]. Communication Design Quarterly Review, 9 (1):39–41, March 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3437000.3437006.

Zhu:2021:MLA

[248] Junzhe Zhu, Elizabeth Wickes, and John R. Gallagher. A machine learning algorithm for sorting online comments via topic modeling. Communication Design Quarterly Review, 9 (2):4–14, July 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453460.3453462.

Butts:2021:RHG

[249] Jimmy Butts and Josephine Walwema.

Rhetorical hedonism and gray genres. Communication Design Quarterly Review, 9(2):15–26, July 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

org/doi/10.1145/3453460.3453461.

Mogull:2021:TCM

[250] Scott A. Mogull. Technical content marketing along the technology adoption lifecycle: experience report. Communication Design Quarterly Review, 9 (2):27–35, July 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3453460.3453463.

Bivens:2021:UHC

[251] Kristin Marie Bivens and Candice A.

Welhausen. Using a hybrid card sortingaffinity diagramming method to teach content analysis: experience report. Communication Design Quarterly Review, 9(3):4–13, September 2021. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3468859.
3468860.

Wright:2021:OTS

[252] David Wright, Daniel B. Shank, and Thomas Yarbrough. Outcomes of training in smart home technology adoption: a living laboratory study. Communication Design Quarterly Review, 9(3): 14–26, September 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3468859.3468861.

Myers:2021:RSV

[253] Angela Myers. Rewriting sexual violence prevention: a comparative rhetorical analysis of online prevention courses in the United States and New Zealand. Communication Design Quarterly Review, 9(3):27–36, September 2021. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3468859.
3468862.

Edenfield:2021:UAP

[254] Avery Edenfield, Hailey Judd, Emmalee Fishburn, and Felicia Gallegos. Unlikely allies in preventing sexual misconduct: Student led prevention efforts in a technical communication classroom: experience report. Communication Design Quarterly Review, 9(4): 4–12, December 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3487213.3487214.

Shafer:2021:RAA

[255] Luana Shafer. Review of “Awful archives: Conspiracy theory, rhetoric, and acts of evidence by Jenny Rice,” Rice, J. (2020). The Ohio State University Press. Communication Design Quarterly Review, 9(4):13–14, December 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3487213.3487215.

Tang:2021:RET

[256] Yingying Tang. Review of “Equipping technical communicators for social justice work: Theories, methodologies, and pedagogies,” by Rebecca Walton & Godwin Y. Agboka; Walton, R., & Agboka, G. Y. (Eds.) (2021). Equipping technical communicators for social justice work: Theories, methodologies, and pedagogies. University Press of Colorado. Communication Design Quarterly Review, 9 (4):15–16, December 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

org/doi/10.1145/3487213.3487216.

Hernandez:2021:RLP

[257] Jess Vázquez Hernández. Review by “Literacy and pedagogy in an age of misinformation and disinformation,” Edited by Tara Lockhart, Brenda Glascott, Chris Warnick, Juli Parrish, and Justin Lewis; Lockhart, T., Glascott, B., Warnick, C., Parrish, J., & Lewis, J. (Eds.) (2021). Parlor Press. Communication Design Quarterly Review, 9(4): 17–18, December 2021. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3487213.3487217.

Sparby:2022:IDU

[258] Erika Sparby and Courtney Cox. Investigating disembodied university crisis communications during COVID-19. Communication Design Quarterly Review, 10(1):4–13, March 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3507454.3507455.

Altamirano:2022:ERS

[259] Amanda Altamirano and Sonia H. Stephens. Experience report streamlining complex website design using a content audit selection heuristic. Communication Design Quarterly Review, 10 (1):14–23, March 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3507454.3507456.

Ballentine:2022:DHT

[260] Brian Ballentine. Digital humanities and technical communication pedagogy: a case and a course for crossprogram opportunities. Communication Design Quarterly Review, 10(1): 24–37, March 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3507454.3507457.

Dayley:2022:EDS

[261] Chris Dayley. Ethical deception: student perceptions of diversity in college recruitment materials. Communication Design Quarterly Review, 10 (1):38–50, March 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3507454.3507458.

Hope:2022:RCB

[262] Lacy Hope. Review of “Composition and Big Data, edited by Amanda Licastro and Benjamin Miller,” (2021). University of Pittsburg Press. Communication Design Quarterly Review, 10 (1):51–53, March 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3507454.3507459.

Durazzi:2022:RTM

[263] Allison Durazzi. Review of “Type Matters: The Rhetoricity of Letterforms edited by Christopher Scott Wyatt and Dànielle Nicole DeVoss,” (2018). Parlor Press. Communication Design Quarterly Review, 10(1): 54–56, March 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3507454.3507460.

Wilkes:2022:RRO

[264] Lydia Wilkes. Review of “Rhet Ops: Rhetoric and Information Warfare edited by Jim Ridolfo and William Hart-Davidson,” (2019). University of Pittsburgh Press. Communication Design Quarterly Review, 10(1):57–59, March 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3507454.3507461.

Frith:2022:ICD

[265] Jordan Frith and Sarah Read. Introduction: communication and design infrastructures. Communication Design Quarterly Review, 10(2):4–9, July 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3507857.
3507858.

Ranade:2022:ISU

[266] Nupoor Ranade and Jason Swarts. Infrastructural support of users’ mediated potential. Communication Design Quarterly Review, 10(2):10–21, July 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3507857.
3507859.

Sherrill:2022:AIP

[267] John T. Sherrill and Michael J. Salvo. Automated infrastructures: participation’s changing role in postindustrial work. Communication Design Quarterly Review, 10(2):22–31, July 2022. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3507857.
3507860.

McMullin:2022:BED

[268] Michelle McMullin, Hadi Riad Banat, Shelton Weech, and Bradley Dilger. Building ethical distributed teams through sustained attention to infrastructure. Communication Design Quarterly Review, 10(2):32–43, July 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3507857.
3507861.

Davis:2022:WIF

[269] Katlynne Davis, Danielle Mollie Stambler, Jessica Lynn Campbell, Daniel L. Hocutt, Ann Hill Duin, and Isabel Pedersen. Writing infrastructure with the fabric of digital life platform. Communication Design Quarterly Review, 10 (2):44–56, July 2022. CODEN ????

ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3507857.3507862.

York:2022:AHC

[270] Eric J. York. Alternate histories and conflicting futures: git version control as software development infrastructure. Communication Design Quarterly Review, 10(2):57–65, July 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

org/doi/10.1145/3507857.3507863.

Lerma:2022:RLC

[271] Corina Lerma. Review of “Literacy as Conversation: Learning Networks in Urban and Rural Communities by Eli Goldblatt and David A. Jolliffe” Goldblatt, E., & Jolliffe, D. A. (2020). University Of Pittsburgh Press. Communication Design Quarterly Review, 10 (2):66–68, July 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3507857.3507864.

Pellegrini:2022:RDT

[272] Mason Pellegrini. Review of “Design Thinking in Technical Communication: Solving Problems through Making and Collaboration by Jason C. K. Tham” Tham, J. C. K. (2021). Routledge. Communication Design Quarterly Review, 10 (2):69–71, July 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3507857.3507865.

Read:2022:IWI

[273] Sarah Read and Jordan Frith. Introduction: writing infrastructure. Communication Design Quarterly Review, 10 (3):5–9, September 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3507870.3507871.

Itchuaqiyaq:2022:CPS

[274] Cana Uluak Itchuaqiyaq and Jordan Frith. Citational practices as a site of resistance and radical pedagogy: positioning the multiply marginalized and underrepresented (MMU) scholar database as an infrastructural intervention. Communication Design Quarterly Review, 10 (3):10–19, September 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3507870.3507872.

Gilbert:2022:TPI

[275] Carrie Anne Gilbert. The textprivileging infrastructures of academic journals. Communication Design Quarterly Review, 10(3):20–21, September 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

                                        URL https://dl.acm.org/doi/10.
                                        1145/3507870.3507873.

Comi:2022:IMS

[276] Dana Comi. “It must be a system thing”: information infrastructure genres as sites of inequity. Communication Design Quarterly Review, 10(3): 22–32, September 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3507870.3507874.

Weber:2022:MIN

[277] Ryan Weber. Making infrastructure into nature: how documents embed themselves into the bodies of oysters. Communication Design Quarterly Review, 10 (3):33–45, September 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

org/doi/10.1145/3507870.3507875.

Adams:2022:TIR

[278] Jonathan Adams. A theory of infrastructural rhetoric. Communication Design Quarterly Review, 10(3): 46–55, September 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3507870.3507876.

Rouge:2022:USA

[279] Mary Le Rouge, Clancy Ratliff, and Donnie Johnson Sackey. Using situational analysis to reimagine infrastructure. Communication Design Quarterly Review, 10(3):56–66, September 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3507870.
3507877.

Caravella:2022:SER

[280] Elizabeth Caravella, Rich Shivener, and Nanditha Narayanamoorthy. Surveying the effects of remote communication & collaboration practices on game developers amid a pandemic. Communication Design Quarterly Review, 10(4): 5–15, December 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3531210.3531211.

Greene:2022:EDA

[281] Jacob Greene. Ethical design approaches for workplace augmented reality. Communication Design Quarterly Review, 10(4):16–26, December 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3531210.
3531212.

Clay:2022:EPR

[282] Michael Clay, Jennifer Smith-Mayo, and Bridie McGreavy. Embodied participation: (re)Situating bodies in collaborative research. Communication Design Quarterly Review, 10(4):27– 39, December 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3531210.3531213.

Friedman:2022:RBM

[283] Malaka Friedman. Review of “Beyond the Makerspace: Making and Relational Rhetorics” by Ann Shivers-McNair (2021). University of Michigan Press. Communication Design Quarterly Review, 10(4):40–41, December 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3531210.
3531214.

Mathis:2022:RVM

[284] Wesley Mathis. Review of “Vibrant Matter: a Political Ecology of Things” by Jane Bennett, (2010). Duke University Press. Communication Design Quarterly Review, 10(4): 42–43, December 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3531210.3531216.

Neal:2022:RDM

[285] D’Arcee Charington Neal. Review of “Dislike-Minded: Media, Audiences, and the Dynamics of Taste” by Jonathan Gray, (2021). New York. Communication Design Quarterly Review, 10(4):44–45, December 2022. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

org/doi/10.1145/3531210.3531215.

Swacha:2023:CCP

[286] Kathryn Yankura Swacha. The Coping with COVID Project: Participatory public health communication. Communication Design Quarterly Review, 11 (1):4–18, March 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3563890.3563891.

Balghare:2023:EHC

[287] Akshata J. Balghare. Exploring healthcare communication gaps between US universities and their international students: a technical communication approach. Communication Design Quarterly Review, 11(1):19–31, March 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3563890.
3563892.

Clem:2023:QNR

[288] Sam Clem and Beth Buyserie. Questioning neoliberal rhetorics of wellness: Designing programmatic interventions to better support graduate instructor wellbeing. Communication Design Quarterly Review, 11(1): 32–41, March 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3563890.3563893.

Carter:2023:CSC

[289] Daniel Carter. Constructing structured content on WordPress: Emerging paradigms in Web content management.

Communication Design Quarterly Review, 11(1):42–52, March 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3563890.3563894.

DeGenaro:2023:RPP

[290] Anthony DeGenaro. Review of “The Profession and Practice of Technical Communication” by Yvonne Cleary (2022). Routledge. Communication Design Quarterly Review, 11(1):53–54, March 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3563890.3563896.

Liu:2023:RED

[291] Meng-Hsien Neal Liu. Review of “Everyday Dirty Work: Invisibility, Communication, and Immigrant Labor” by Wilfredo Alvarez, (2022). The Ohio State University Press. Communication Design Quarterly Review, 11(1): 55–57, March 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3563890.3563895.

Amidon:2023:CER

[292] Timothy R. Amidon, Kristen R. Moore, and Michele Simmons. Community engaged researchers and designers: How we work and what we need. Communication Design Quarterly Review, 11(2):5–9, July 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3592356.3592357.

Grant:2023:DPI

[293] Carrie Grant and Dorian Walker. Designing public identity: Finding voice in coalitional technical writing with black-led organizations. Communication Design Quarterly Review, 11(2): 10–17, July 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592356.3592358.

Wertz:2023:SSB

[294] Olivia M. Wertz, Kandi Workman, and Erin Brock Carlson. Seeking out the stakeholders: Building coalitions to address cultural (in)equity through artsbased, community-engaged research. Communication Design Quarterly Review, 11(2):18–27, July 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

org/doi/10.1145/3592356.3592359.

Novotny:2023:CDC

[295] Maria Novotny, Gina Davis, Maya Grobel, and Jennifer Vesbit. Communitydriven concepts to support TPC coalition building in a post-Roe world. Communication Design Quarterly Review, 11 (2):28–37, July 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592356.3592360.

Allison:2023:MGS

[296] Lydia Allison, Christopher Maggio, Salma Kalim, and Megan Schoettler. Making graduate student CER practices visible: Navigating the double-binds of identities, space, and time. Communication Design Quarterly Review, 11 (2):38–43, July 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592356.3592361.

Hartline:2023:HLS

[297] Megan Faver Hartline. The hidden labor of sustaining community partnerships. Communication Design Quarterly Review, 11(2):44–49, July 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

                                        org/doi/10.1145/3592356.3592362.

Amidon:2023:ISI

[298] Timothy R. Amidon, Ann Blakeslee, Erin Brock Carlson, Lehua Ledbetter, Kristen R. Moore, Emma Rose, and Michele Simmons. Introduction to the second issue: a conversation about community-engaged research. Communication Design Quarterly Review, 11 (3):5–11, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3592367.3592368.

Itchuaqiyaq:2023:DCE

[299] Cana Uluak Itchuaqiyaq, Chris A. Lindgren, and Corina Qaagraq Kramer. Decolonizing community-engaged research: Designing CER with cultural humility as a foundational value. Communication Design Quarterly Review, 11(3): 12–20, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3592367.3592369.

Ledbetter:2023:BPW

[300] Lehua Ledbetter and Alexandria Neelis. Beyond policy: What plants and communities can teach us about sustainable changemaking. Communication Design Quarterly Review, 11(3):21– 27, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592367.3592370.

Lee:2023:CEU

[301] Soyeon Lee, Heather Noel Turner, and Emma J. Rose. Community-engaged user experience pedagogy: Stories, emergent strategy, and possibilities. Communication Design Quarterly Review, 11(3):28–41, September 2023. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3592367.
3592371.

Blakeslee:2023:SCB

[302] Ann M. Blakeslee, Kristine M. Gatchel, David Boeving, and Brent Miller. Story of a community-based writing resource — and a call to engage. Communication Design Quarterly Review, 11(3): 42–53, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592367.3592372.

Cameron:2023:ADN

[303] Shanna Cameron. Amplifying diverse narratives of social support in online health design. Communication Design Quarterly Review, 11(3): 54–66, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592367.3592373.

Moore:2023:TDC

[304] Kristen R. Moore and Erica M. Stone.

Tracing the development and circulation of a tool for coalitional change. Communication Design Quarterly Review, 11 (3):67–72, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

org/doi/10.1145/3592367.3592374.

Lantz:2023:RVE

[305] Susan Jennings Lantz. Review of “Violent Exceptions: Children’s Human Rights and Humanitarian Rhetorics by Wendy S. Hesford,” The Ohio State University Press. Communication Design Quarterly Review, 11(3): 73–74, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592367.3592375.

Das:2023:RUE

[306] Meghalee Das. Review of “User Experience as Innovative Academic Practice by Kate Crane and Kelli Cargile Cook,” The WAC Clearinghouse; University Press of Colorado. https://doi.org/ 10.37514/TPC-B.2022.1367. Communication Design Quarterly Review, 11(3): 75–77, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592367.3592376.

Coulter:2023:RTS

[307] Andi Coulter. Review of “Tuning in to Soundwriting by Kyle D. Stedman, Courtney S. Danforth, & Michael J. Faris,” Stedman, K. D., Danforth, C. S., & Faris, M. J. (Eds.). (2021). enculturation/Intermezzo. http://intermezzo.enculturation.net/14stedman-et-al/index.html. Communication Design Quarterly Review, 11(3): 78–79, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592367.3592377.

Hines:2023:RWC

[308] Jasara Hines. Review of “Writing in the Clouds: Inventing and Composing in Internetworked Writing Spaces by John Logie,” (2021). Parlor Press. Communication Design Quarterly Review, 11(3): 80–81, September 2023. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3592367.3617935.

Carradini:2024:FIR

[309] Stephen Carradini and Barbara Carradini. Fernweh interdisciplinary research visualizer: a data visualization tool for interdisciplinary research scoping. Communication Design Quarterly Review, 12 (1):1–10, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3627691.3627692.

Mussack:2024:CSV

[310] Brigitte Mussack and Jason Tham.

Collaboration as a shared value: Instructor and student perceptions of collaborative learning in online business writing courses. Communication Design Quarterly Review, 12(1): 11–25, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3627691.3627693.

Stevens:2024:PAT

[311] Hannah L. Stevens. Publicly available, transparent, and explicit: an analysis of academic publishing policy and procedure documents. Communication Design Quarterly Review, 12 (1):26–36, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3627691.3627694.

Tetu:2024:DAW

[312] Imari Cheyne Tetu, Shannon Kelly, Jun Fu, Caitlin K. Kirby, Scott Schopieray, and Stephen Thomas. Developing asynchronous workshop models for professional development. Communication Design Quarterly Review, 12(1): 37–43, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3627691.3627695.

Eccles:2024:IME

[313] Kathryn Eccles, Laura Herman, Caterina Moruzzi, and Maggie Mustaklem. Introducing the method of exhibit-based research. Communication Design Quarterly Review, 12(1): 44–49, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3627691.3627696.

Lauren:2024:PID

[314] Benjamin Lauren. The political impact of the default of GenAI. Communication Design Quarterly Review, 12 (1):50–52, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3627691.3627697.

Gray:2024:RSE

[315] Kat M. Gray. Review of Salt of the earth: Rhetoric, preservation and white supremacy by James Chase Sanchez, Sanchez, J. C. (2021). Conference on College Composition and Communication, NCTE Press. Communication Design Quarterly Review, 12(1):

53–54, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3627691.3627698.

Puntasecca:2024:RPM

[316] Tina Puntasecca. Review of Privacy matters: Conversations about surveillance in and beyond the classroom by Estee Beck and Les Hutchinson Campos,” Beck E, & Hutchinson Campos, L. (Eds.). (2020). Utah State University Press. Communication Design Quarterly Review, 12(1): 55–57, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3627691.3627699.

Reamer:2024:REE

[317] David Reamer. Review of Embodied Environmental Risk in Technical Communication by Samuel Stinson and Mary Le Rouge,” Stinson, S., & Le Rouge, M. (Eds.). (2022). Routledge. Communication Design Quarterly Review, 12 (1):58–59, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3627691.3627700.

Virtue:2024:RUC

[318] Drew Virtue. Review of Update culture and the afterlife of digital writing by John R. Gallagher,” Gallagher, J. R. (2019). Utah State University Press. Communication Design Quarterly Review, 12(1):60–61, March 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https://dl.acm.

org/doi/10.1145/3627691.3627701.

Stambler:2024:TDL

[319] Danielle Mollie Stambler, Saveena (Chakrika) Veeramoothoo, and Katlynne Davis. Toward digital life: Embracing, complicating, and reconceptualizing digital literacy in communication design. Communication Design Quarterly Review, 12(2):5–10, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3655727.3655728.

Chen:2024:PDL

[320] Chen Chen. The post-digital life of transnational activists: Develop a tactical technological literacy. Communication Design Quarterly Review, 12 (2):11–21, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3655727.3655729.

Buck:2024:RMD

[321] Amber Buck. Redrawing the maps: Digital literacy practices of grassroots activists. Communication Design Quarterly Review, 12(2):22–32, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/3655727.
                                        3655730.

Heilig:2024:AAE

[322] Leah Heilig, Ally Overbay, Madison Jones, and Taylor Roberts. Augmenting for accessible environments: Layering deep mapping, deep accessibility, and community literacy. Communication Design Quarterly Review, 12 (2):33–43, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3655727.3655731.

Gallagher:2024:ASA

[323] Philip B. Gallagher and Marci J.

Gallagher. Accessible sound: Aural information literacy for technical communication design. Communication Design Quarterly Review, 12(2): 44–52, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3655727.3655732.

Karr:2024:DGL

[324] Danielle Feldman Karr, Steve Holmes, Jared S. Colton, and Josephine Walwema. The digital “Good life”: The limits of applying an ethics of care to a company “Running with scissors”. Communication Design Quarterly Review, 12 (2):53–63, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3655727.3655733.

Tofteland-Trampe:2024:WTZ

[325] Rachel Tofteland-Trampe. Writing in the “Twilight Zone” and lessons for inclusive design. Communication Design Quarterly Review, 12(2): 64–73, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3655727.3655734.

Swartz:2024:BLT

[326] Haley Swartz. Biodigital literacy through intimate data: User perceptions of FemTech and pelvic floor training devices. Communication Design Quarterly Review, 12(2):74–85, June 2024. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3655727.
3655735.

Woods:2024:RDP

[327] Charles Woods and Gavin P. Johnson. (Re)Designing privacy literacy in the age of generative AI. Communication Design Quarterly Review, 12 (2):86–97, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3655727.3655736.

Dammeyer:2024:SSD

[328] Sarah Dammeyer. Special section dedicated to Dr. Halcyon Lawrence: “Please continue this good work!”: a letter to Dr. Halcyon Lawrence from a brief friend. Communication Design Quarterly Review, 12(2): 98–99, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3655727.3655737.

Hutter:2024:HDH

[329] Liz Hutter. Honoring Dr. Halcyon Lawrence’s legacy in the technical communication classroom. Communication Design Quarterly Review, 12(2): 100–101, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3655727.3655738.

Mainaly:2024:GLL

[330] Shiva Hari Mainaly. A glimpse of Lawrence’s legacy: From “Siri Discipline” to disciplining artificial intelligence. Communication Design Quarterly Review, 12(2):102–104, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/3655727.
                                        3655739.

Neefe:2024:DHL

[331] Lauren Neefe. Dr. Halcyon Lawrence:

a resounding legacy. Communication Design Quarterly Review, 12(2): 105–107, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3655727.3655740.

Ojedele-Adejumo:2024:DHL

[332] Temitope Ojedele-Adejumo. Dr. Halcyon Lawrence’s “Siri Disciplines”: Examining accented English and pedagogical implications of biased technologies through an African diasporic lens. Communication Design Quarterly Review, 12 (2):108–110, June 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3655727.3655741.

Rose:2024:UPS

[333] Emma J. Rose and Heather Noel Turner.

UX pedagogy: Stories and practices from the technical and professional communication classroom. Communication Design Quarterly Review, 12(3): 1–4, September 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3658422.3658423.

Rea:2024:CEE

[334] Ashley Rea and Akshata Balghare.

Cultivating empathic engineering design through UX pedagogy: Challenges and insights. Communication Design Quarterly Review, 12(3): 5–19, September 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3658422.3658429.

Smith:2024:SPD

[335] Allegra W. Smith and Courtney D. Ray. 17 students, 1 project: Design thinking pedagogy for a large-scale UX community/classroom partnership. Communication Design Quarterly Review, 12(3): 20–30, September 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3658422.3658430.

Hunter:2024:BMO

[336] Paul Thompson Hunter. Balancing methodological openness and control in TPC-UX pedagogy. Communication Design Quarterly Review, 12(3): 31–40, September 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3658422.3658426.

Donelson:2024:VVT

[337] Jerrice Renita Donelson. Voices from the void: Teaching user experience as racial storytelling in TPC. Communication Design Quarterly Review, 12(3): 41–48, September 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3658422.3658432.

Patcha:2024:ERO

[338] Sidouane Patcha and Sarah Read. An experience report on the opportunities and challenges of a community-engaged user experience (CEUX) pedagogy for a masters-level course. Communication Design Quarterly Review, 12(3): 49–58, September 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

                                        tronic). URL https://dl.acm.org/
                                        doi/10.1145/3658422.3658428.

Johnson:2024:TLD

[339] Emily K. Johnson. Teaching liberatory design. Communication Design Quarterly Review, 12(3):59–70, September 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3658422.3658427.

Strantz:2024:IUB

[340] Adam Strantz. Interactive UX: Building and testing for accessibility with design systems. Communication Design Quarterly Review, 12(3):71–79, September 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3658422.3658431.

Atherton:2024:BET

[341] Rachel Atherton and Patricia Brackin.

Building empathy through classroom and community integration in a multidisciplinary engineering design program. Communication Design Quarterly Review, 12(3):80–87, September 2024. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic). URL https:

//dl.acm.org/doi/10.1145/3658422.
3658424.

Gomes:2024:LLB

[342] Mathew Gomes. Localizing laborbased contract grading for a communityengaged UX course. Communication Design Quarterly Review, 12(3):88– 100, September 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (elec-

tronic). URL https://dl.acm.org/
doi/10.1145/3658422.3658425.

Lancaster:2024:ETA

[343] Amber Lancaster and Carie S. Tucker King. Empowerment through authorship inclusivity: Toward more equitable and socially just citation practices. Communication Design Quarterly Review, 12(4):1–15, December 2024. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Lambrecht:2024:HVM

[344] Kathryn Lambrecht. Heat vulnerability mapping: Designing visual tools that effectively communicate risk. Communication Design Quarterly Review, 12(4): 16–26, December 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Bakke:2024:AIL

[345] Abigail Bakke and Leonard DiBono. Are academia and industry listening to each other? A citation analysis of UX research methods resources. Communication Design Quarterly Review, 12(4): 27–38, December 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Liddle:2024:CCD

[346] Daniel Liddle. Commemoration and context: The death counter graphics of COVID-19. Communication Design Quarterly Review, 12(4):39–53, December 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

VanWinkle:2024:BLV

[347] Kevin Van Winkle. Between the lines: Visual euphemism in technical and professional communication visuals. Communication Design Quarterly Review, 12 (4):54–63, December 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Johnson:2024:RDJ

[348] Nathan R. Johnson. Review of “Data Justice and the Right to the City” by Morgan Currie, Jeremy Knox and Callum McGreggor (Eds.) (2022). Edinburgh University Press. Communication Design Quarterly Review, 12(4): 64–65, December 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Mainaly:2024:RBA

[349] Shiva Hari Mainaly. Review of “Book Anatomy: Body Politics and Materiality of Indigenous Book History” by Amy Gore, (2023). University of Massachusetts Press. Communication Design Quarterly Review, 12(4):66–68, December 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Tang:2024:RAV

[350] Yingying Tang. Review of “Amplifying Voices in UX: Balancing Design and User Needs in Technical Communication” by Amber L. Lancaster and Carie S. T. King (Eds.), SUNY Press. Communication Design Quarterly Review, 12 (4):69–70, December 2024. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Wang:2024:REW

[351] Hua Wang. Review of “Engineering Words: Communicating Clearly in the Workplace” by Sharon Burton and Bonni Graham Gonzalez, XML Press. Communication Design Quarterly Review, 12(4):71–72, December 2024. CO-DEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Johnson:2025:DAT

[352] Meredith A. Johnson and Lauren E. Cagle. Designing accessible tables: When technical communication competencies come into conflict. Communication Design Quarterly Review, 13(1):1–14, March 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Xu:2025:UDV

[353] Qian Xu and Ke Jiang. UX design vs. UI design: Understanding U.S. employers’ expectations through semantic analysis of job descriptions. Communication Design Quarterly Review, 13(1):15–30, March 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Friess:2025:SCW

[354] Erin Friess. The second coder was a robot: Generative AI tools in establishing intercoder reliability. Communication Design Quarterly Review, 13 (1):31–39, March 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Craska:2025:DEE

[355] Heather D. Craska, Amélie Y. Davis, and W. Michele Simmons. Designing for engagement: Evaluating perception of quick-response (QR) codes in informal environmental education and outreach materials. Communication Design Quarterly Review, 13(1):40–50, March 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Weedon:2025:ASK

[356] Scott Weedon and Charlie Rioux. Articulating science: Knowledge translation as a methodology for scientific and technical communication research. Communication Design Quarterly Review, 13 (1):51–58, March 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Gresbrink:2025:RVT

[357] Emily Gresbrink. Review of “Voice and Tone Strategy: Connecting with People through Content” by John Caldwell,, Laguna Hills, CA: XML Press. Communication Design Quarterly Review, 13 (1):59–60, March 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Pandya:2025:RUE

[358] Manushri K. Pandya. Review of “User Experience as Innovative Academic Practice” by Kate Crane and Kelli Cargile Cook,” University Press of Colorado. Communication Design Quarterly Review, 13(1):61–63, March 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Tudor:2025:RIG

[359] Benjamin Tudor. Review of “Interrogating Gendered Pathologies” by Erin A. Frost and Michelle F. Eble (Eds.), University Press of Colorado. Communication Design Quarterly Review, 13 (1):64–65, March 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Littlefield:2025:TTC

[360] Jamie Littlefield. From tactical technical communication to infrastructural writing: The role of user enfranchisement in a rogue street design manual. Communication Design Quarterly Review, 13(2):1–11, June 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Ross:2025:AEZ

[361] Derek G. Ross. Aberrance as expansion: Zines, acceptability ethics, and radical communication in technical & professional communication. Communication Design Quarterly Review, 13(2):12–21, June 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Sindelar:2025:AER

[362] Kailan Sindelar, Juliette Guido, Kaylee Klosterman, and Brooke Wooten. The AR elephant in the room: a method for user experience research in AR photography apps. Communication Design Quarterly Review, 13(2):22–31, June 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Andrews:2025:RFH

[363] M. L. Andrews. Review of “Fitter, Happier: The Eugenic Strain in Twentieth-Century Cancer Rhetoric by Lois Peters Agnew,” Agnew, L. P. (2024). The University of Alabama Press. Communication Design Quarterly Review, 13(2):32– 33, June 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Green:2025:RUW

[364] Barbara C. G. Green. Review of “UX Writing: Designing User-Centered Content by Jason C. K. Tham, Tharon Howard, and Gustav Verhulsdonck,” (1st ed.). Routledge. https://doi.org/10.4324/978-1-003-27441-4. Communication Design Quarterly Review, 13(2):34–35, June 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).

Warner:2025:RUE

[365] Chip Warner. Review of “User Experience Research and Usability of Health Information Technology by Jessica Lynn Campbell, PhD”, CRC Press. Communication Design Quarterly Review, 13 (2):36–38, June 2025. CODEN ???? ISSN 2166-1200 (print), 2166-1642 (electronic).